### v1

# Adaptive Hyper-Graph Convolution Network for Skeleton-based Human Action Recognition with Virtual Connections

Abstract
The shared topology of human skeletons motivated the recent investigation of graph convolutional network (GCN) solutions for action recognition. However, the existing GCNs rely on the binary connection of two neighbouring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures. In this paper we address this oversight and explore the merits of a hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices. In particular, our Hyper-GCN adaptively optimises multi-scale hyper-graphs during training, revealing the action-driven multi-vertex relations. Besides, virtual connections are often designed to support efficient feature aggregation, implicitly extending the spectrum of dependencies within the skeleton. By injecting virtual connections into hyper-graphs, the semantic clues of diverse action categories can be highlighted. The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets, demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods. Specifically, we outperform the existing solutions on NTU-120, achieving 90.2% and 91.4% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
Index Terms:
Action recognition, Video Understanding, Computer Vision, Graph Learning.
I Introduction
Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR, and AR [1, 2, 3, 4, 5, 6, 7]. In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms. Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension and higher efficiency. Moreover, it is robust to illumination changes and scene variations. For these reasons, the adoption of this structural data is very popular in skeleton-based action recognition [8, 9, 10, 11, 12, 13, 14].
Figure 1: Illustration of our Hyper-GCN. In (a), dotted lines represent the virtual connections. Each coloured part containing multiple joints represents a hyper-graph with hyper-edges. (b) provides an intuitive comparison between normal graph convolution and hyper-graph convolution operations with the same connection degree.
To facilitate skeleton-based action recognition, both Recurrent Neural Networks (RNNs) [8, 9, 10, 11] and Convolutional Neural Networks (CNNs) [15, 16, 17] have been well explored. However, RNNs themselves cannot depict the intrinsic skeleton topology. The learned filters of CNNs, on the other hand, neglect the spatio-temporal structure of the skeleton. Drawing on these observations, recent studies have focused on how to model the skeleton topology directly.
In principle, the physical topology of human joints and bones can be consistently represented by a graph. Accordingly, the graph convolutional network (GCN) is typically introduced to achieve the aggregation of feature information conveyed by skeleton joints [18, 13, 19]. In general, two neighbouring joints can pass messages through their shared bone. In graph terms, this corresponds to the exchange of information between two vertices along their connecting edge. Besides the physical skeleton topology, attempts have been made to explore implicit relationships among joints [20, 13, 12, 19], suggesting different variants of skeleton topology. In principle, these approaches assume a binary connection between each connected vertex pair. Mathematically, the constructed topology in the adjacency matrix is represented by a normal graph. However, human actions are jointly defined by several joints. Hence, human actions encompass not only binary relations between vertex pairs but also multi-joint relationships. For instance, the action primitive starting running is manifest in the human raising the left hand while the right leg steps forward, as shown in Figure 1 (a). The binary connections are not sufficient to capture the synergistic interaction of multiple joints. This strongly argues for constructing feature aggregation paths involving multiple vertices.
Accordingly, we propose to construct a hyper-graph to depict the skeleton topology and take advantage of the outstanding performance of hyper-graph analysis techniques  [21, 22, 23, 24]. The hyper-graph topology involves multiple node connections, rather than binary connections of a normal graph. As shown in Figure 1 (a), in a hyper-graph, a hyper-edge can link more than two vertices, which has the capacity to represent complex collaborative relations among human joints. As one hyper-edge associates multiple vertices, a single hyper-graph convolution enables aggregating all the features along the hyper-edges. An illustration is provided in Figure 1 (b), where the normal graph and hyper-graph are presented to demonstrate their differences in passing the information conveyed by the vertex features.
Conceptually, for a fair comparison, we typically set each hyper-edge to contain only 2 vertices, which is the same as the normal graph. In the normal graph convolution, after 2-layer aggregation, the information of the green vertex is spread to 2 other vertices. In contrast, in a hyper-graph convolution, the information of the green vertex spreads to 4 vertices including itself, creating an extended receptive field. Theoretically, by applying a hyper-graph for modelling joint aggregation on a skeleton, improved efficiency can be obtained during learning the action semantics.
Besides its information aggregation mode, the capacity of an action recognition system is also limited by its input features. In general, the number of skeleton joints is fixed in existing benchmarks, e.g., 25 for NTU-120 [25]. The entire recognition process relies on the interactions among the involved joints. In the domain of artistic puppetry, the history of driving actions, such as Shadow Play  [^1] https://en.wikipedia.org/wiki/Shadow_play and Marionette  [^2] https://en.wikipedia.org/wiki/Marionette goes back more than 2000 years. This kind of art form provides an inspiration for involving additional ’hyper joints’ which can drive or facilitate a better communication between existing joints. The underlying spirit is to alleviate the pressure of the real joints in storing and transferring semantics. Jointly with the hyper joints, real joints can focus more on storing neighbouring joint features and transferring the global information to the hyper joints. Interestingly, the class token in existing Transformers can also be considered as a hyper-token  [26, 27, 28, 29]. As shown in Figure 1, the skeleton is described as a marionette, where the actions are ”controlled” by connecting hyper joints to the real joints. This suggests that the hyper joints are not only able to capture the representation information of human action, but also reveal the implicit information between physically connected joints as hints for recognition.
By endowing a hyper-graph with hyper joints, virtual connections are created to perform comprehensive hyper-graph convolutions. We construct Hyper-GCN based on the above design principles. Extensive experiments on 3 datasets, NTU RGB+D 60, NTU RGB+D 120, and NW-UCLA, are conducted for evaluation. The results validate the merits of our proposed Hyper-GCN. The main contributions are as follows:
  • We propose to represent human skeleton topology by a hyper-graph. Compared with the normal graph, the efficiency of feature interaction is improved via multi-vertex aggregation.
  • Virtual hyper joints are injected to connect the physical joints, enhancing the model capacity to capture the global semantics.
  • We propose an adaptive hyper-graph convolution network (Hyper-GCN) for skeleton-based human action recognition. On all three public datasets used for experimentation, Hyper-GCN achieves the SOTA performance.
Figure 2: The framework of our Hyper-GCN. Part (a) represents the Multi-Scale Hyper-GC (MS-HGC) module, where channels are divided into 8 branches. Part (b) represents the process of constructing an adaptive hyper-graph. Part (c) represents the architecture of Hyper-GCN. $\mathbf{F}_{h}$ represents the hyper-joints which are learnable parameters. $\hat{A}_{p}$ represents the physical topology. $HA$ represents The fused topology of the adaptive hyper-graph and the human physical structure.
II Related Work
II-A Graph Topology for Action Recognition
A graph can represent a human skeleton, preserving the joint relationships via predefined edges. To aggregate semantics, GCNs have been well studied in recent years [30, 31, 32, 33, 34, 35, 36, 37, 38]. Regarding modelling the spatio-temporal relevance, ST-GCN [18] proposes to represent the topology with an adjacency matrix of 3 subsets. Similarly, 2s-AGCN [20], InfoGCN [12], and DS-GCN [14] use a self-attentive mechanism to learn the topology from joint pairs. Besides the intuitive spatio-temporal dimensions, CTR-GCN [13] proposes to learn the channel topology using subspace mappings. All the above methods construct the topology from joint pairs, which rely only on the binary relation between two vertices. In their implementation stage, the adjacency matrix is used to represent the topology. In this case, high-order information among joints is not taken into consideration, neglecting the collaborative power among multiple joints. Though hyper-graph is considered to construct topology using multivariate joint relationships, current [39, 40] solutions manually set the hyper-graphs, which greatly relies on human experience, sacrificing the adaptability of graph learning.
II-B Feature Settings for Action Recognition
It has been observed that the input features play the most essential role in delivering high recognition accuracy [41, 42, 43, 36, 44]. Drawing on this, Graph2Net [45] proposes to extract local and global spatial features jointly. Similarly, CTR-GCN [13] uses multi-scale temporal convolution to extract temporal features. While HD-GCN [19] focuses on learning the feature aggregation among different edge sets. To extend the perception field, STC-Net [46] uses the dilated kernels for graph convolution to capture the features. After all, existing GCNs [18, 20, 12, 14, 47] receive the input features from real skeleton joints. However, each skeleton joint acts as an information carrier during forward passing, which is required to deliver both local context and global semantics. Given a fixed representative capacity (number of channels), we believe it is necessary to involve additional virtual joints to balance the pressure of storing local and global information.
III Approach
III-A Preliminaries
In general, the input features in graph convolution are represented by $\mathbf{{F}_{in}}\in\mathbb{R}^{C\times T\times V}$, where $C$ represents the number of channels in the feature maps. Given the topology of a human skeleton, we usually define the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of joints and $\mathcal{E}$ represents the set of edges between joints. For the set of edges $\mathcal{E}$, it is formulated as an adjacent matrix $A\in\mathbb{R}^{N\times N}$, where $N$ represents the number of human joints. The normalised adjacency matrix is represented by $\hat{A}\in\mathbb{R}^{N\times N}$. The normalisation operation is formulated as follows:
$\hat{A}=\Lambda^{-\frac{1}{2}}\ A\Lambda^{-\frac{1}{2}},$ (1)
where $\Lambda$ represents the diagonal matrix stored with the degree of every joint. The entire normal graph convolution process can be formulated as:
$\mathbf{F_{out}}=\sigma(\hat{A}\mathbf{F_{in}}W),$ (2)
where $W\in\mathbb{R}^{C\times C^{\prime}}$ is the learnable parameters, representing the feature transformation patterns in the feature space. $\sigma$ denotes the non-linear activation function ReLU.
III-B Hyper-graph
Here, we use $\mathcal{G_{H}}=(\mathcal{V_{H}},\mathcal{E_{H}},\mathcal{W_{H}})$ to define the spatial hyper-graph [24] with human skeleton. $\mathcal{V_{H}}$ and $\mathcal{E_{H}}$ follow similar definitions in the normal graph, which represent the set of joints and the set of hyper-edges. In addition, we introduce $\mathcal{W_{H}}$ to represent the weights of each hyper-edge. Since a hyper-edge contains multiple joints, the corresponding topology can no longer be simply represented by an adjacency matrix. Therefore, we introduce the incidence matrix $H\in\mathbb{R}^{N\times M}$ to describe the topology of each joint in the hyper-graph. $N$ represents the number of joints and $M$ represents the number of hyper-edges. Given $v\in\mathcal{V_{H}}$ and $e\in\mathcal{E_{H}}$, the values of incidence matrix can be determined by:
$h(e,v)=\begin{cases}1,&v\in e,\\ 0,&v\notin e\end{cases}.$ (3)
Similar to the normal graph convolution, it also needs to normalise the hyper-graph to modulate the aggregated features. The degree of hyper-graph consists of the degree of joints and the degree of hyper-edges. The degree of joints is represented by the sum of weights of all joints contained in each hyper-edge. Given $v\in\mathcal{V_{H}}$, it can be described as follows:
$d(v)\ =\sum_{e\in\mathcal{E_{H}}}{w(e)h(v,e)},$ (4)
where we use a diagonal matrix $D_{v}\in\mathbb{R}^{N\times N}$ to represent the degree of joints. $W\in\mathbb{R}^{M\times M}$ represents the weight matrix of hyper-edges, which is formulated as a diagonal matrix. The degree of hyper-edges represents the sum of the number of joints contained in each hyper-edge. Given $e\in\mathcal{E_{H}}$, it can be described as follows:
$d(e)\ =\sum_{v\in\mathcal{V_{H}}}{h(v,e)},$ (5)
where we use a diagonal matrix $D_{e}\in\mathbb{R}^{M\times M}$ to represent the degree of hyper-edges. Therefore, the normalisation of a hyper-graph is formulated as follows:
$\hat{H}=D_{v}^{-1}HWD_{e}^{-1}H^{T},$ (6)
where $\hat{H}\in\mathbb{R}^{N\times N}$ represents the normalised incidence matrix for hyper-graph convolution.
III-C Adaptive Hyper-graph Construction Module
How to construct the hyper-graph is critical to better understanding human actions. As shown in Figure 2 (b), we design an Adaptive Hyper-graph Construction Module (AHC-Module). Given the feature $\mathbf{{F}^{{}^{\prime}}_{in}}\in\mathbb{R}^{C\times T\times V}$, we first perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction. This can efficiently decouple the temporal and spatial clues. After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}^{{}^{\prime}}_{in}}\in\mathbb{R}^{C\times V}$. To adjust the scale of the input feature channels, we adopt a layer normalisation operation on the channel dimension. After that, we use two individual mapping functions to embed the features into the subspace for constructing the hyper-graph. This design enables preserving the original features to reflect their spatial relevance. The embedding operation can be formulated as follows:
$P_{\textrm{emb}}$ $=\Phi(\textrm{LN}(\textrm{Pooling}_{\textrm{T}}(\mathbf{F^{{}^{\prime}}_{in}}))),$ (7)
$W_{\textrm{emb}}$ $=\textrm{Tanh}(\Psi(\textrm{LN}(\textrm{Pooling}_{\textrm{T}}(\mathbf{F^{{}^{\prime}}_{in}})))),$ (8)
where $P_{\textrm{emb}}$ represents the relative position of the joints in the embedding space, and $W_{\textrm{emb}}$ represents the weights of hyper-edges. $\textrm{Pooling}_{\textrm{T}}$ denotes performing pooling in the temporal dimension. LN denotes the layer normalisation along the channel dimension. $\Phi\in\mathbb{R}^{C\times C_{h}},\Psi\in\mathbb{R}^{C\times 1}$ are the mapping functions of position and weight, which are set learnable parameters. $C_{h}$ represents the channels of the mapping subspace. Tanh denotes the activation function to obtain the weight of each hyper-edge, limiting the values with the range of $[-1,1]$.
Figure 3: Illustration of the hyper-graph construction process with the proposed AHC-Module. $N$ represents the number of joints.
In the embedding space, our AHC-Module utilises the Euclidean distance to measure the joint difference. For the mapping of hyper-edges and vertices, the reason we apply only one mapping function instead of two separate ones is to ensure that vertices and hyper-edges are within the same semantic space. Define the set of joints $P_{\textrm{emb}}=\{v_{0}\in\mathbb{R}^{C_{h}},v_{1}\in\mathbb{R}^{C_{h}},...,v_{N}\in\mathbb{R}^{C_{h}}\}$, given $v_{i},v_{j}\in P_{\textrm{emb}}$, each element $m_{i,j}$ in the distance matrix $M\in\mathbb{R}^{N\times N}$ can be obtained as follows:
$dis_{i,j}=dis_{j,i}=||v_{i}-v_{j}||^{2}_{2}.$ (9)
In principle, we need to transform the distance matrix into a probability hyper-grap0h, which determines whether the specific joints are contained with the same hyper-edges. In order to guarantee that the model is trainable, we use the softmax operation to assign the probabilities. However, directly using the softmax would inevitably result in each joint belonging to all hyper-edges to some extent. Since this outcome is undesirable, we introduce constraints on each joint to limit the number of hyper-edges it can belong to, preventing excessive connections. We retain only the $K$ nearest hyper-edges represented by each joint in the semantic space for probability estimation. Given the row vector $\mathbf{m}_{i}$ in the distance matrix $M$, which represents the distances between joint $i$ and the other joints, we select the indices of the minimum $K$ joints from $\mathbf{m}_{i}$ and add them to the set $set^{K}_{i}$. Based on $set^{K}_{i}$, $h_{i,j}$ in the incidence matrix $H\in\mathbb{R}^{N\times N}$ of hyper-graph, can be calculated as follows:
$h_{i,j}$ =$=\begin{cases}\frac{e^{-m_{i,j}}}{\sum_{k\in set^{K}_{i}}{e^{-m_{i,k}}}},&j\in set^{K}_{i},\\ 0,&j\notin set^{K}_{i}\end{cases}$ (10)
The shape of our incidence matrix is consistent with the adjacency matrix of the normal graph. While our topology with hyper-graph contains more high-order information. The entire process is illustrated in the Figure 3.
III-D Multi Scale Hyper-graph Convolution
As our AHC-Module can depict the topology into a hyper-graph at different scales by adjusting the value of $K$. To accommodate the semantic information reflected by these varying scales, we further propose the Multi-Scale Hyper-graph Convolution (MS-HGC), as shown in Figure 2 (a). We use different branches to independently perform hyper-graph convolution on the topological information represented by the multi-scale hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN.
Besides the hyper-graph, we incorporate the physical topology to emphasise the natural physical relations of a human being’s skeleton. To achieve this objective, existing methods [18, 20, 13, 19] divide the physical topology of the human body into 3 subsets. They are represented by $S=\{s_{id},s_{cf},s_{cp}\}$, where $s_{id}$, $s_{cf}$ and $s_{cp}$ denote the identity, centrifugal, and centripetal joint subsets. In our MS-HGC, to ensure that the integrated physical topology in each branch remains complete, we merge them into a single set. We divide the input feature into 8 separate branches along the channel dimension, which are processed by 8 MS-HGC parallelly, delivering 8 multi-scale hyper-graphs. 8 branches are aggregated by channel concatenation after the hyper-graph convolutions, as shown in Figure 2 (a). The operation of MS-HGC can be formulated as follows:
$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha_{i}\cdot\hat{H_{i}})\mathbf{F^{[(i-1)\times C^{{}^{\prime}}+1,...,i\times C^{{}^{\prime}}]}_{in}}W_{i}},$ (11)
[…]
III-E Virtual Connections
It is worth emphasising that incorporating learnable joints among different samples is essential for enhancing the model capacity, as these learnable joints capture generalised features of human actions. This not only enriches the semantic information but also facilitates easier interaction connections among real joints. Therefore, we introduce the hyper-joints which are to participate in the hyper-graph convolution as shown in Figure 2 (a). The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C^{{}^{\prime}}\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C^{{}^{\prime}}\times T\times V_{h}}$. Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data. Furthermore, we set independent hyper-joints at each layer of Hyper-GCN, aiming to harmonise features at different depths. Typically, these hyper-joints are involved in spatial hyper-graph convolution rather than in temporal convolution.
To diversify these hyper-joints, we propose the Divergence Loss for hyper-joints optimisation to mitigate their homogenisation. We adopt a cosine matrix $C\in\mathbb{R}^{V_{h}\times V_{h}}$ to measure the differences between hyper-joints, which can be formulated as follows:
$C=\frac{\text{ReLU}(\mathbf{F_{h}F^{T}_{h}})}{||\mathbf{F_{h}}||^{2}}-\mathbf{I},$ (12)
where $\mathbf{I}$ represents the identity matrix which is used to emphasise the irrelevance between each learned hyper-joint pair.
Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer. The calculation can be formulated as:
$\mathcal{L}_{h}(C)$ $=\frac{\sum^{V_{h}}_{i=1}\sum^{V_{h}}_{j=1}c_{i,j}}{(V_{h}-1)^{2}},\quad c_{i,j}\in C,$ (13)
$\mathcal{L}$ $=\mathcal{L}_{CE}+\frac{1}{L}\sum^{L}_{l=1}L_{h}(C_{l}),$ (14)
where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer. Additionally, to preserve the original physical topology, we use the AHC-Module solely to learn the topology between feature $\mathbf{F_{h}}$ and feature $\mathbf{F_{p}}$.
III-F Entire Architecture
Hyper-GCN consists of an embedding layer, 9 spatial-temporal convolution layers, a global average pooling, and an FC layer, as shown in Figure 2 (c).
The embedding layer is utilised to map individual joints into a high-dimensional feature space for subsequent adaptive hyper-graph convolution. To preserve the position awareness of the embedded features, we introduce the position embedding in the embedding layer inspired by InfoGCN [12].
The 9 layers of Hyper-GCN are categorised into 3 stages. Each stage consists of 3 layers. Each layer consists of Multi-Scale Hyper-graph convolution (MS-HGC) and Multi-Scale Temporal convolution (MS-TC) [13]. In stage 1, the number of channels is set as 128. Then it increases to 256 for stage 2 and stage 3.
Dense connections are introduced to integrate deep and shallow features in each stage, which can smooth the obtained distribution against the hyper-vertices interaction at each layer.
Methods Publication Modalities GFLOPs Params (M) NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
ST-GCN [18] AAAI 2018 J+B - - 81.5 88.3 70.7 73.2 -
2s-AGCN [20] CVPR 2019 J+B - - 88.5 95.1 82.5 84.2 -
DGNN [32] CVPR 2019 J+B+JM+BM - - 89.9 96.1 - - -
SGN [30] CVPR 2020 J+B+JM+BM - - 89.0 94.5 79.2 81.5 92.5
DC-GCN+ADG [33] ECCV 2020 J+B+JM+BM 1.83 4.9 90.8 96.6 86.5 88.1 95.3
DDGCN [31] ECCV 2020 J+B+JM+BM - - 91.1 97.1 - - -
MS-G3D [42] CVPR 2020 J+B+JM+BM 5.22 2.8 91.5 96.2 86.9 88.4 -
MST-GCN [34] AAAI 2021 J+B+JM+BM - 12.0 91.5 96.6 87.5 88.8 -
CTR-GCN [13] ICCV 2021 J+B+JM+BM 1.97 1.5 92.4 96.4 88.9 90.6 96.5
EfficientGCN-B4 [43] TPAMI 2022 J+B+JM+BM 15.2 2.0 91.7 95.7 88.3 89.1 -
InfoGCN* [12] CVPR 2022 J+B+J’+B’ 1.84 1.6 92.7 96.9 89.4 90.7 96.6
FR Head [48] CVPR 2023 J+B+JM+BM - 2.0 92.8 96.8 89.5 90.9 96.8
HD-GCN* [19] ICCV 2023 J+B+J’+B’ 1.77 1.7 93.0 97.0 89.8 91.2 96.9
DS-GCN [14] AAAI 2024 J+B+JM+BM - - 93.1 97.5 89.2 90.3 -
Hyper-GCN (w/o ensemble) - J 1.62 1.0 91.1 95.3 87.0 88.4 94.8
Hyper-GCN (2-ensemble) - J+B 1.62 1.0 92.9 96.3 89.8 90.9 96.2
Hyper-GCN (4-ensemble) - J+B+JM+BM 1.62 1.0 93.1 96.7 90.2 91.4 96.6
TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets. For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods. J, B, JB, and JM represent the joint, bone, joint motion, and bone motion. In addition, for InfoGCN, HD-GCN [12, 19], we use the results from 4-streams according to the original papers for a fair comparison. GFLOPs and Params (M) all represent the single-stream.
IV Evaluation
In this section, we conduct extensive experiments on standard benchmarking datasets to evaluate our Hyper-GCN, in the task of skeleton-based human action recognition.
IV-A Datasets
NTU-RGB+D 60 NTU-RGB+D 60 [11] is a large dataset widely used in skeleton-based human action recognition. It consists of 56,880 different samples, categorised into 60 classes, obtained from the directed performances of 40 different actors. There are two 2 evaluation benchmarks. (1) Cross-Subject (X-Sub): For the 40 actors, 20 are used for training and 20 for validation. (2) Cross-view (X-View): For 3 views, 2 are used for training, and 1 for validation.
NTU-RGB+D 120 NTU-RGB+D 120 [25] is am extention of NTU 60, which introduce 57,367 new action samples. It consists of 114,480 different samples, categorised into 120 classes instead of 60 classes. The number of actors increase to 106. It also corresponds yo 2 benchmarks. (1) Cross-Subject (X-Sub): For the 106 actors, 53 are used for training and 53 for validation. (2) Cross-Setup (X-Set): For the 32 setups, the sample with even setup IDs are used for training, and the odd setup IDs for validation.
Northwestern-UCLA. The Northwestern-UCLA (NW-UCLA) dataset [49] contains 1494 video clips, which is categorised into 10 classes. It is obtained through 3 cameras with different camera views and is performed by 10 actors. As the proposed benchmark, for 3 camera views, 2 camera views are used for training, and the remaining 1 is used for validation.
IV-B Implementation Details
Our training and evaluation stages are on a single RTX 3090 GPU. In terms of the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004. Our implementation uses label smooth cross-entropy loss with the Divergence Loss we proposed. We set a total of 140 epochs with the start 5 warm-up epochs. The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120. The training batch size is set to 64 for NTU RGB+D and NTU RGB+D 120, and 16 for Northwestern-UCLA.
IV-C Comparison with the State-of-the-Art
Multi-stream ensembles proposed in  [50] have been proven effective by most of the existing state-of-the-art methods. We also use a 4-streams ensemble to evaluate our model.
The detailed results are reported in Table I. To the best of our knowledge, a comparison with SOTA shows that our proposed Hyper-GCN achieves the SOTA level on each dataset. Especially, we outperform the involved competitors on NTU-RGB+D 120. This demonstrates the effectiveness and merits of Hyper-GCN.
IV-D Ablation Study
We conduct ablation experiments and report visualisations to analyse the effectiveness of each component of our design. All the ablation study is evaluated on the benchmark of NTU-RGB+D 120, using the joint modality. The Baseline for comparison is set by replacing the MS-HGC with vanilla GC, and removing the hyper-joints and dense connections in the backbone.
Figure 4: Top-1 accuracy in selected classes. The $K$-scale represents the $K$-scale hyper-graph in MS-HGC. The data is displayed starting from 75%.
$K$-scale
NTU RGB+D 120
X-Sub (%)
Baseline - 85.1
Single scale 3 86.4 ($\uparrow$ 1.3)
5 86.5 ($\uparrow$ 1.4)
7 86.5 ($\uparrow$ 1.4)
9 86.3 ($\uparrow$ 1.2)
Multi scale [0,2,3,4,5,6,7,8] 86.6 ($\uparrow$ 1.5)
[2,3,4,5,6,7,8,9] 86.7 ($\uparrow$ 1.6)
[3,4,5,6,7,8,9,10] 86.6 ($\uparrow$ 1.5)
TABLE II: Comparison of the hyper-parameter $K$ in terms of adaptive hyper-graph construction. [$k_{1}$,$k_{2}$,…,$k_{8}$] represents the $K$-scale hyper-graph of 8 branches in MS-HGC. The bold font represents the best results.
Multi-Scale Hyper-graph: As we mentioned, Multi-Scale Hyper-graph Convolution (MS-HGC) consists of an 8-branch parallel architecture. In particular, each branch corresponds to a hyper-graph at a different scale. The hyper-graphs at different scales represent the maximum number of hyper-edges each joint is contained in. They reflect the varying degrees of aggregation in the constructed topology. The different scales of hyper-graphs are determined by the hyper-parameter $K$ that we define. Therefore, to explore how to set the hyper-parameter $K$ for the branches in MS-HGC, we design the ablation experiments to investigate the impact of different scales on the model’s recognition ability, as well as to compare the results by defining a single scale.
The results are reported in Table II. When $K$ equals 1, each hyper-edge contains only one joint, which essentially eliminates any interaction between joints. Therefore, we do not set K to 1. Observing the design of the single scale hyper-graph, it becomes clear that increasing $K$ does not always yield better performance. A very large $K$ means that the maximum number of hyper-edges each joint is contained in increases, which in turn causes the number of joints in each hyper-edge to grow. Even though these ”extra” joints may be optimised for minimal weight, they are likely to act as ”noise”, disrupting the information interaction represented by the hyper-edge.
Furthermore, compared to a single-scale design, it is evident that the multi-scale design of MS-HGC outperforms the single scale design. For the complex and diverse characteristics of human action, multi-scale hyper-graphs in the spatial dimension can better capture features at multiple levels. Therefore, we visualise the Top-1 accuracy of selected human action categories as shown in Figure 4. This is especially advantageous for recognising actions with varying degrees of motion. For instance, for actions like ”Brushing teeth” or ”Salutation”, where only local joints are in motion, the smaller-scale hyper-graphs focus more on the local joints. In contrast, for actions like ”Stand up” and ”Jump up”, where the entire body is engaged, a larger-scale hyper-graph can capture information between distant joints. This also inspired us to adopt the AHC-Module in different branches to complementarily learn hyper-graph topology at different scales.
NTU RGB+D 120
X-Sub (%)
GFLOPs Parmas (M)
Baseline 85.1 1.70 1.52
w/ MS-HGC 86.6 ($\uparrow$ 1.5) 1.59 1.03
w/ 1 86.7 ($\uparrow$ 1.6) 1.61 1.03
w/ 2 86.9 ($\uparrow$ 1.8) 1.62 1.04
w/ 3 87.0 ($\uparrow$ 1.9) 1.62 1.05
w/ 4 86.6 ($\uparrow$ 1.5) 1.64 1.05
w/ 5 86.4 ($\uparrow$ 1.3) 1.66 1.06
TABLE III: The comparison of the configurations of hyper joints. w/ $N$ represents the number of hyper joints in each layer. These experiments are conducted in the setting w/ MS-HGC. The bold font represents the best results.
Figure 5: The comparison of the cosine matrices of hyper-joints. The matrix elements range from $[-1,1]$. This experiment is set w/ 3 hyper-joints. Part (a) represents w/o Divergence Loss. Part (b) represents w/ Divergence Loss. The cosine matrix is calculated by Eqn. 12.
Virtual Connections: We conduct the ablation study to analyse the hyper joints with the Divergence Loss. The results are listed in Table III. By comparison, the performance is not monotonically increasing with the number of hyper joints. The case of introducing only 3 hyper joints achieves the best performance. As our hyper joints are learned from a large amount of data, too many hyper joints can introduce redundant and ambiguous clues, weakening the model capacity. Besides, the involvement of hyper joints delivers consistent improvement, while introducing only a marginal increase in terms of model size and complexity.
In addition, we visualised the cosine matrices in certain layers of Hyper-GCN, as shown in Figure 5. Clearly, in the absence of the Divergence Loss, the homogenisation of the learned hyper-joints in several layers is severe. This can limit the ability of hyper-joints to represent the generalised features of human actions. The observation further validates the effectiveness of the Divergence Loss for hyper-joint optimisation, supporting our viewpoint that generalised features with a certain degree of differentiation are needed to participate in hyper-graph convolution.
Method S-1 S-2 S-3 MS-HGC H NTU-RGB+D 120
X-Sub
Baseline 84.8
$\checkmark$ 84.8
$\checkmark$ 84.8
$\checkmark$ 84.9 ($\uparrow$ 0.1)
Hyper-GCN $\checkmark$ $\checkmark$ 84.9 ($\uparrow$ 0.1)
$\checkmark$ $\checkmark$ 85.0 ($\uparrow$ 0.2)
$\checkmark$ $\checkmark$ $\checkmark$ 85.1 ($\uparrow$ 0.3)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 86.6 ($\uparrow$ 2.0)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 87.0 ($\uparrow$ 2.4)
TABLE IV: The ablation study of architecture settings. S-$i$ represents adding dense connection in stage $i$. MS-HGC and H represent w/ MS-HGC and w/ 3 hyper-joints.
Figure 6: t-SNE visualisation of Baseline and Hyper-GCN. Different colours represent the features of different joints.
Architecture of Backbone: We validate the effectiveness of performing dense connections in the backbone. The experimental results are collected in Table IV. The dense connections in the experiments are only added within each stage. It is observed that dense connections in shallow layers do not improve the performance as much as in deep layers. As the layers go deeper, the number of channels gradually increases and the variance of information extracted from each layer becomes steeper. This suggests that the deeper stages require dense connections to modulate the variance between features. For example, the performance introducing dense connections at stage 1 and stage 2 is close to adding the dense connection only at stage 3.
To further validate that Hyper-GCN increases the efficiency of information interaction, we perform t-SNE projections on the output features of each layer, as shown in Figure 6. Notice that the semantic information represented by the last layer features of Hyper-GCN is very convergent. This represents that the semantic information of each joint is adequately conveyed in Hyper-GC. The joints with similar semantic can prevent sacrificing the information after global average pooling.
V Conclusion
In this paper, we propose a Hyper-graph Convolutional Network for skeleton-based human action recognition. To exploit the implicit topology of multivariate synergy between joints, we introduce Multi-Scale Hyper-graph Convolution and virtual connections. We also introduce dense connections to fuse shallow and deep features. We carry out experiments on the dataset NTU-RGB+D 60 & 120, NW-UCLA to validate the effectiveness of Hyper-GCN. The experimental analysis verifies that our design can improve the recognition performance. To the best of our knowledge, Hyper-GCN achieves the SOTA performance on 3 public datasets. The involvement of hyper-graph modelling indeed extends existing GCN-based action recognition paradigms.

### v2

@@
 The shared topology of human skeletons motivated the recent investigation of graph convolutional network (GCN) solutions for action recognition.
-However, the existing GCNs rely on the binary connection of two neighbouring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures.
-In this paper we address this oversight and explore the merits of a hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices.
-In particular, our Hyper-GCN adaptively optimises multi-scale hyper-graphs during training, revealing the action-driven multi-vertex relations.
+However, most of the existing GCNs rely on the binary connection of two neighboring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures.
+Although some studies have attempted to utilize hyper-graphs to represent the topology, they rely on a fixed construction strategy, which limits their adaptivity in uncovering the intricate latent relationships within the action.
+In this paper, we address this oversight and explore the merits of an adaptive hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices.
+In particular, our Hyper-GCN adaptively optimises the hyper-graphs during training, revealing the action-driven multi-vertex relations.
 Besides, virtual connections are often designed to support efficient feature aggregation, implicitly extending the spectrum of dependencies within the skeleton.
 By injecting virtual connections into hyper-graphs, the semantic clues of diverse action categories can be highlighted.
-The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets, demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods.
-Specifically, we outperform the existing solutions on NTU-120, achieving 90.2% and 91.4% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
+The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods.
+Specifically, we outperform the existing solutions on NTU-120, achieving 90.5% and 91.7% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
 Index Terms:
@@
 In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms.
-Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension and higher efficiency.
+Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension, conveying human action with higher efficiency.
 Moreover, it is robust to illumination changes and scene variations.
@@
 Each coloured part containing multiple joints represents a hyper-graph with hyper-edges.
-(b) provides an intuitive comparison between normal graph convolution and hyper-graph convolution operations with the same connection degree.
+(b) provides an intuitive comparison between normal graph convolution and hyper-graph convolution operations with the same connectivity degree.
 To facilitate skeleton-based action recognition, both Recurrent Neural Networks (RNNs) [8, 9, 10, 11] and Convolutional Neural Networks (CNNs) [15, 16, 17] have been well explored.
@@
 In principle, the physical topology of human joints and bones can be consistently represented by a graph.
-Accordingly, the graph convolutional network (GCN) is typically introduced to achieve the aggregation of feature information conveyed by skeleton joints [18, 13, 19].
+Accordingly, the graph convolutional network (GCN) is typically introduced to aggregate feature information conveyed by skeleton joints [18, 19, 20].
+In terms of data representations, skeleton joints are also one-dimensional sequence data, similar to language.
+Therefore, with the rise of recent self-attention techniques [21, 22], several Transformer-based action recognition solutions [23, 24, 25, 26] have been released.
+Transformer considers skeletal joints as tokens and uses attention maps to reflect the topology.
+These methods have achieved excellent performance but sacrifice much higher GFLOPS and parameters than GCN.
 In general, two neighbouring joints can pass messages through their shared bone.
 In graph terms, this corresponds to the exchange of information between two vertices along their connecting edge.
-Besides the physical skeleton topology, attempts have been made to explore implicit relationships among joints [20, 13, 12, 19], suggesting different variants of skeleton topology.
+Besides the physical skeleton topology, attempts have been made to explore the implicit relationships among joints [27, 13, 12, 19], suggesting different variants of skeleton topology.
 In principle, these approaches assume a binary connection between each connected vertex pair.
@@
 The hyper-graph topology involves multiple node connections, rather than binary connections of a normal graph.
-As shown in Figure 1 (a), in a hyper-graph, a hyper-edge can link more than two vertices, which has the capacity to represent complex collaborative relations among human joints.
+As shown in Figure 1 (a), in a hyper-graph, a hyper-edge can link more than two vertices.
+This linking has the capacity to represent complex collaborative relations among human joints.
 As one hyper-edge associates multiple vertices, a single hyper-graph convolution enables aggregating all the features along the hyper-edges.
 An illustration is provided in Figure 1 (b), where the normal graph and hyper-graph are presented to demonstrate their differences in passing the information conveyed by the vertex features.
-Conceptually, for a fair comparison, we typically set each hyper-edge to contain only 2 vertices, which is the same as the normal graph.
 In the normal graph convolution, after 2-layer aggregation, the information of the green vertex is spread to 2 other vertices.
-In contrast, in a hyper-graph convolution, the information of the green vertex spreads to 4 vertices including itself, creating an extended receptive field.
-Theoretically, by applying a hyper-graph for modelling joint aggregation on a skeleton, improved efficiency can be obtained during learning the action semantics.
-Besides its information aggregation mode, the capacity of an action recognition system is also limited by its input features.
+In contrast, in a hyper-graph convolution, the information of the green vertex spreads to all the vertices, creating an extended receptive field.
+Theoretically, by modelling a skeleton using a hyper-graph for the aggregation of joint information, improved efficiency can be obtained during learning the action semantics.
+However, the key to achieving enhanced information aggregation by hyper-graph convolution lies in the conformity of the hyper-graph structure to the real topology of human action.
+Several studies [32, 33] attempt to adopt a hyper-graph to model the skeleton topology, mostly based on superficial prior knowledge and an unfounded definition of the structure of hyper-graph.
+Distinguishing our approach from the existing hyper-graph-based methods, we design an adaptive solution for constructing the hyper-graph with virtual connections.
+It can construct a Non-uniform hyper-graph specific to each individual action type.
+Besides the information aggregation issue, the capacity of an action recognition system is also constrained by its input features.
 In general, the number of skeleton joints is fixed in existing benchmarks, e.g., 25 for NTU-120 [34].
-The entire recognition process relies on the interactions among the involved joints.
+The entire recognition process relies on a successful capture of the interactions among the skeleton joints.
 In the domain of artistic puppetry, the history of driving actions, such as Shadow Play  [^1] https://en.wikipedia.org/wiki/Shadow_play and Marionette  [^2] https://en.wikipedia.org/wiki/Marionette goes back more than 2000 years.
 This kind of art form provides an inspiration for involving additional ’hyper joints’ which can drive or facilitate a better communication between existing joints.
-The underlying spirit is to alleviate the pressure of the real joints in storing and transferring semantics.
-Jointly with the hyper joints, real joints can focus more on storing neighbouring joint features and transferring the global information to the hyper joints.
-Interestingly, the class token in existing Transformers can also be considered as a hyper-token  [26, 27, 28, 29].
+The underlying spirit is to alleviate the pressure on the real joints to store and transfer complex semantics.
+Jointly with the hyper joints, real joints can focus more on storing neighbouring joint features and relegating the task of transferring the global information to the hyper joints.
+Interestingly, the class token in existing Transformers can also be considered as a hyper-token  [21, 22].
 As shown in Figure 1, the skeleton is described as a marionette, where the actions are ”controlled” by connecting hyper joints to the real joints.
 This suggests that the hyper joints are not only able to capture the representation information of human action, but also reveal the implicit information between physically connected joints as hints for recognition.
-By endowing a hyper-graph with hyper joints, virtual connections are created to perform comprehensive hyper-graph convolutions.
+By endowing an adaptive non-uniform hyper-graph with hyper joints, virtual connections are created to perform comprehensive hyper-graph convolutions.
 We construct Hyper-GCN based on the above design principles.
-Extensive experiments on 3 datasets, NTU RGB+D 60, NTU RGB+D 120, and NW-UCLA, are conducted for evaluation.
+Extensive experiments on 3 datasets, NTU RGB+D 60 [11], NTU RGB+D 120 [34], and NW-UCLA [35], are conducted for evaluation.
 The results validate the merits of our proposed Hyper-GCN.
 The main contributions are as follows:
-  • We propose to represent human skeleton topology by a hyper-graph. Compared with the normal graph, the efficiency of feature interaction is improved via multi-vertex aggregation.
-  • Virtual hyper joints are injected to connect the physical joints, enhancing the model capacity to capture the global semantics.
-  • We propose an adaptive hyper-graph convolution network (Hyper-GCN) for skeleton-based human action recognition.
-On all three public datasets used for experimentation, Hyper-GCN achieves the SOTA performance.
+  • An adaptive non-uniform hyper-graph to represent the human skeleton topology. Compared with a fixed hyper-graph, the constructed topology is more action-specific, thus boosting the discrimination.
+  • The injection of virtual hyper joints, enriching the connectivity of the physical joints, from a global semantic perspective.
+  • As the processing architecture, we propose a hyper-graph convolution network (Hyper-GCN).
+The SOTA performance achieved on three public datasets demonstrates the merits of the Hyper-GCN and virtual connection designs.
 Figure 2: The framework of our Hyper-GCN.
@@
 A graph can represent a human skeleton, preserving the joint relationships via predefined edges.
-To aggregate semantics, GCNs have been well studied in recent years [30, 31, 32, 33, 34, 35, 36, 37, 38].
-Regarding modelling the spatio-temporal relevance, ST-GCN [18] proposes to represent the topology with an adjacency matrix of 3 subsets.
+To aggregate semantics, GCNs [36, 37, 38, 39, 40, 41, 42, 20] and Transformers  [23, 43, 25, 24, 44] have been well studied.
+For GCN-based approaches, ST-GCN [18] proposes to represent the topology with an adjacency matrix of 3 subsets by modeling the spatio-temporal relevance.
 Similarly, 2s-AGCN [27], InfoGCN [12], and DS-GCN [14] use a self-attentive mechanism to learn the topology from joint pairs.
-Besides the intuitive spatio-temporal dimensions, CTR-GCN [13] proposes to learn the channel topology using subspace mappings.
-All the above methods construct the topology from joint pairs, which rely only on the binary relation between two vertices.
-In their implementation stage, the adjacency matrix is used to represent the topology.
+Besides the intuitive spatio-temporal dimensions, CTR-GCN [13] proposes to learn the topology for channels to refine the skeleton.
+For Transformer-based solutions, IIP-Transformer [26] adopt a self-attention mechanism to establish the intra-part and the inter-part joints.
+IGFormer [45] learns the topology between persons at both semantic and distance levels.
+STTFormer [46] proposed structure can capture the correlation between joints in consecutive frames.
+All the above two paradigms construct the topology from joint pairs, which rely only on the binary relation between two vertices.
+In their implementation stage, the adjacency matrix or attention map is used to represent the topology.
 In this case, high-order information among joints is not taken into consideration, neglecting the collaborative power among multiple joints.
 Though hyper-graph is considered to construct topology using multivariate joint relationships, current [33, 32] solutions manually set the hyper-graphs, which greatly relies on human experience, sacrificing the adaptability of graph learning.
-II-B Feature Settings for Action Recognition
-It has been observed that the input features play the most essential role in delivering high recognition accuracy [41, 42, 43, 36, 44].
+II-B Feature Configurations for Action Recognition
+It has been observed that the input features play the most essential role in delivering high recognition accuracy [47, 48, 49, 42, 50, 26, 45, 46] in GCNs and Transformers.
 Drawing on this, Graph2Net [51] proposes to extract local and global spatial features jointly.
-Similarly, CTR-GCN [13] uses multi-scale temporal convolution to extract temporal features.
-While HD-GCN [19] focuses on learning the feature aggregation among different edge sets.
+CTR-GCN [13] uses multi-scale temporal convolution to extract temporal features.
+While HD-GCN [19] introduce the hierarchical edge convolution to key edge features.
 To extend the perception field, STC-Net [52] uses the dilated kernels for graph convolution to capture the features.
-After all, existing GCNs [18, 20, 12, 14, 47] receive the input features from real skeleton joints.
+Similarly, IG-Transformer [45] aggregate the features of two persons as one to mine the semantic information.
+STST [43] split the coordinate features, semantic features and temporal features into three kinds of tokens.
+TSTE [24] extracting and merging the spatial and temporal features as input to transformer encoders.
+After all, existing methods [18, 27, 12, 14, 53, 25, 23] receive the input features from real skeleton joints.
 However, each skeleton joint acts as an information carrier during forward passing, which is required to deliver both local context and global semantics.
@@
 where $\Lambda$ represents the diagonal matrix stored with the degree of every joint. The entire normal graph convolution process can be formulated as:
-$\mathbf{F_{out}}=\sigma(\hat{A}\mathbf{F_{in}}W),$ (2)
-where $W\in\mathbb{R}^{C\times C^{\prime}}$ is the learnable parameters, representing the feature transformation patterns in the feature space.
+$\mathbf{F_{out}}=\sigma(\hat{A}\mathbf{F_{in}}P),$ (2)
+where $P\in\mathbb{R}^{C\times C^{\prime}}$ is the learnable parameters, representing the feature transformation patterns in the feature space.
 $\sigma$ denotes the non-linear activation function ReLU.
@@
 where we use a diagonal matrix $D_{e}\in\mathbb{R}^{M\times M}$ to represent the degree of hyper-edges. Therefore, the normalisation of a hyper-graph is formulated as follows:
-$\hat{H}=D_{v}^{-1}HWD_{e}^{-1}H^{T},$ (6)
+$\hat{H}=D_{v}^{-\frac{1}{2}}HWD_{e}^{-1}H^{T}D_{v}^{-\frac{1}{2}},$ (6)
 where $\hat{H}\in\mathbb{R}^{N\times N}$ represents the normalised incidence matrix for hyper-graph convolution.
-III-C Adaptive Hyper-graph Construction Module
-How to construct the hyper-graph is critical to better understanding human actions.
-As shown in Figure 2 (b), we design an Adaptive Hyper-graph Construction Module (AHC-Module).
-Given the feature $\mathbf{{F}^{{}^{\prime}}_{in}}\in\mathbb{R}^{C\times T\times V}$, we first perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction.
-This can efficiently decouple the temporal and spatial clues.
-After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}^{{}^{\prime}}_{in}}\in\mathbb{R}^{C\times V}$.
-To adjust the scale of the input feature channels, we adopt a layer normalisation operation on the channel dimension.
-After that, we use two individual mapping functions to embed the features into the subspace for constructing the hyper-graph.
+Figure 3: Instance of the incidence matrix construction by A-NHG with hyper-parameter $K=3$.
+$N$ represents the number of joints.
+(a) represents the hyper-edge sets are null.
+(b) represents that measure the distance between 1-st joint and other joints.
+(c) represents that retrain $K$-nearest hyper-edges which contain 1-st joint.
+(d) represents the final state.
+III-C Adaptive Non-uniform Hyper-graph
+Constructing the hyper-graph incidence matrix is critical to understanding human actions. As shown in Figure 2 (b), we design an Adaptive Non-uniform Hyper-graph (A-NHG) Construction.
+Given the general feature $\mathbf{X}\in\mathbb{R}^{N\times C}$ of spatial skeleton joints, $N$ represent the number of joints, $C$ represent the channels.
+Our A-NHG construct the hyper-edge with each joint as the center of mass, and a total of $N$ hyper-edges are constructed.
+A-NHG utilises the Euclidean distance to measure the joint difference.
+We use one mapping function $\Phi\in\mathbb{R}^{C\times C_{h}}$ to embed the features into the subspace for constructing the hyper-graph.
+$C_{h}$ represents the hidden channels of the mapping subspace.
 This design enables preserving the original features to reflect their spatial relevance.
-The embedding operation can be formulated as follows:
-$P_{\textrm{emb}}$ $=\Phi(\textrm{LN}(\textrm{Pooling}_{\textrm{T}}(\mathbf{F^{{}^{\prime}}_{in}}))),$ (7)
-$W_{\textrm{emb}}$ $=\textrm{Tanh}(\Psi(\textrm{LN}(\textrm{Pooling}_{\textrm{T}}(\mathbf{F^{{}^{\prime}}_{in}})))),$ (8)
-where $P_{\textrm{emb}}$ represents the relative position of the joints in the embedding space, and $W_{\textrm{emb}}$ represents the weights of hyper-edges.
-$\textrm{Pooling}_{\textrm{T}}$ denotes performing pooling in the temporal dimension.
-LN denotes the layer normalisation along the channel dimension.
-$\Phi\in\mathbb{R}^{C\times C_{h}},\Psi\in\mathbb{R}^{C\times 1}$ are the mapping functions of position and weight, which are set learnable parameters.
-$C_{h}$ represents the channels of the mapping subspace.
-Tanh denotes the activation function to obtain the weight of each hyper-edge, limiting the values with the range of $[-1,1]$.
-Figure 3: Illustration of the hyper-graph construction process with the proposed AHC-Module. $N$ represents the number of joints.
-In the embedding space, our AHC-Module utilises the Euclidean distance to measure the joint difference.
-For the mapping of hyper-edges and vertices, the reason we apply only one mapping function instead of two separate ones is to ensure that vertices and hyper-edges are within the same semantic space.
-Define the set of joints $P_{\textrm{emb}}=\{v_{0}\in\mathbb{R}^{C_{h}},v_{1}\in\mathbb{R}^{C_{h}},...,v_{N}\in\mathbb{R}^{C_{h}}\}$, given $v_{i},v_{j}\in P_{\textrm{emb}}$, each element $m_{i,j}$ in the distance matrix $M\in\mathbb{R}^{N\times N}$ can be obtained as follows:
-$dis_{i,j}=dis_{j,i}=||v_{i}-v_{j}||^{2}_{2}.$ (9)
-In principle, we need to transform the distance matrix into a probability hyper-grap0h, which determines whether the specific joints are contained with the same hyper-edges.
+Based on $X_{H}\in\mathbb{R}^{N\times C_{h}}$, we define the set of joints $\mathcal{V_{H}}=\{v_{1}\in\mathbb{R}^{C_{h}},v_{2}\in\mathbb{R}^{C_{h}},...,v_{N}\in\mathbb{R}^{C_{h}}\}$.
+Given $v_{i},v_{j}\in\mathcal{V_{H}}$, each element $m_{i,j}$ in the distance matrix $M\in\mathbb{R}^{N\times N}$ can be obtained as follows:
+$m_{i,j}=m_{j,i}=||v_{i}-v_{j}||^{2}_{2}.$ (7)
+In principle, we need to transform the distance matrix into a probability incidence matrix, which determines whether the specific joints are contained with the same hyper-edges.
 In order to guarantee that the model is trainable, we use the softmax operation to assign the probabilities.
@@
 Since this outcome is undesirable, we introduce constraints on each joint to limit the number of hyper-edges it can belong to, preventing excessive connections.
-We retain only the $K$ nearest hyper-edges represented by each joint in the semantic space for probability estimation.
-Given the row vector $\mathbf{m}_{i}$ in the distance matrix $M$, which represents the distances between joint $i$ and the other joints, we select the indices of the minimum $K$ joints from $\mathbf{m}_{i}$ and add them to the set $set^{K}_{i}$.
-Based on $set^{K}_{i}$, $h_{i,j}$ in the incidence matrix $H\in\mathbb{R}^{N\times N}$ of hyper-graph, can be calculated as follows:
-$h_{i,j}$ =$=\begin{cases}\frac{e^{-m_{i,j}}}{\sum_{k\in set^{K}_{i}}{e^{-m_{i,k}}}},&j\in set^{K}_{i},\\ 0,&j\notin set^{K}_{i}\end{cases}$ (10)
-The shape of our incidence matrix is consistent with the adjacency matrix of the normal graph.
-While our topology with hyper-graph contains more high-order information.
-The entire process is illustrated in the Figure 3.
-III-D Multi Scale Hyper-graph Convolution
-As our AHC-Module can depict the topology into a hyper-graph at different scales by adjusting the value of $K$.
-To accommodate the semantic information reflected by these varying scales, we further propose the Multi-Scale Hyper-graph Convolution (MS-HGC), as shown in Figure 2 (a).
-We use different branches to independently perform hyper-graph convolution on the topological information represented by the multi-scale hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN.
+For one joint, we retain only the $K$ nearest hyper-edges represented by joints in the semantic space for probability estimation.
+And $K$ is the hyper-parameter of A-NHG.
+Given the row vector $\mathbf{m}_{i}$ in the distance matrix $M$, which represents the distances between joint $i$ and the other joints.
+We select the indices of the minimum $K$ joints from $\mathbf{m}_{i}$ and add them to the set $set_{i}$.
+Based on $set_{i}$, $h_{i,j}$ in the incidence matrix $H\in\mathbb{R}^{N\times N}$, can be calculated as follows:
+$h_{i,j}$ =$=\begin{cases}\frac{\text{exp}(-m_{i,j})}{\sum_{k\in set_{i}}{\text{exp}(-m_{i,k})}},&j\in set_{i},\\ 0,&j\notin set_{i}\end{cases}$ (8)
+In contrast to uniform hyper-graph (every hyper-edge contains the same number of joints), our approach limits the maximum number of every joint that can be contained by hyper-edges.
+Obviously, the number of joints contained in each hyper-edge constructed in this way is non-fixed, allowing the hyper-edges to capture more differentiated associations.
+The entire process for example with $K$ = 3 is illustrated in the Figure 3.
+III-D Multi-head Hyper-graph Convolution
+To accommodate the semantic information reflected by different group channels, we further propose the Multi-head Hyper-graph Convolution (M-HGC), as shown in Figure 2 (a).
+We use separate branches to independently perform hyper-graph convolution on the topologies represented by the multi-head hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN.
+For instance, the feature $\mathbf{F_{in}\in\mathbb{R}^{C_{in}\times T\times V}}$ is divided into $8$ separate branches along the channel dimension, which are processed by $8$ M-HGC parallelly, delivering $8$ separate hyper-graphs.
+Next, we perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction.
+This can efficiently decouple the temporal and spatial clues.
+After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}_{in}}\in\mathbb{R}^{C_{in}\times V}$.
+The hyper-graph is depicted by the incidence matrix and weight matrix.
+So, we use a separate mapping function for each head to embed the features into the subspace for constructing the hyper-graph.
+This design enables preserving the original features to reflect their spatial relevance.
+After that, we introduce the A-NHG to obtain the incidence matrix.
+For the weight matrix, we adopt the Channel Attention [54].
+The hyper-graph obtained operation can be formulated as follows:
+$H$ $=\uplus^{8}_{i=1}\text{A-NHG}(\mathbf{\bar{F}^{k}_{in}}),$ (9)
+$W$ $=\sigma^{2}(\Psi^{2}(\uplus^{8}_{i=1}\sigma^{1}(\Psi^{1}_{i}(\mathbf{\bar{F}^{k}_{in}})))),$ (10)
+[…]
 Besides the hyper-graph, we incorporate the physical topology to emphasise the natural physical relations of a human being’s skeleton.
@@
 They are represented by $S=\{s_{id},s_{cf},s_{cp}\}$, where $s_{id}$, $s_{cf}$ and $s_{cp}$ denote the identity, centrifugal, and centripetal joint subsets.
-In our MS-HGC, to ensure that the integrated physical topology in each branch remains complete, we merge them into a single set.
-We divide the input feature into 8 separate branches along the channel dimension, which are processed by 8 MS-HGC parallelly, delivering 8 multi-scale hyper-graphs.
-8 branches are aggregated by channel concatenation after the hyper-graph convolutions, as shown in Figure 2 (a).
-The operation of MS-HGC can be formulated as follows:
-$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha_{i}\cdot\hat{H_{i}})\mathbf{F^{[(i-1)\times C^{{}^{\prime}}+1,...,i\times C^{{}^{\prime}}]}_{in}}W_{i}},$ (11)
-[…]
+In our M-HGC, to ensure that the integrated physical topology in each head remains complete, we merge them into a single set.
+8 heads are aggregated by channel concatenation after the hyper-graph convolutions, as shown in Figure 2 (a).
+The operation of M-HGC can be formulated as follows:
+$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha\cdot\hat{H_{i}})\mathbf{F^{k}_{in}}P_{i}},$ (11)
+$\hat{A}$ and $\hat{H}$ denote the normalized physical adjacency matrix and normalized incidence matrix.
+$P$ is the learnable weight parameter for feature transform.
+$\alpha$ is a learnable parameter for the topology fusion of each head.
+Category Methods Publication Modalities NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
+X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
+GCN-based MST-GCN [40] AAAI21 J+B+JM+BM 91.5 96.6 87.5 88.8 -
+CTR-GCN [13] ICCV21 J+B+JM+BM 92.4 96.4 88.9 90.6 96.5
+EfficientGCN-B4 [49] TPAMI22 J+B+JM+BM 91.7 95.7 88.3 89.1 -
+InfoGCN [12] CVPR22 J+B+JM+BM 92.7 96.9 89.4 90.7 96.6
+FR Head [55] CVPR23 J+B+JM+BM 92.8 96.8 89.5 90.9 96.8
+HD-GCN* [19] ICCV23 J+B+J’+B’ 93.0 97.0 89.8 91.2 96.9
+DS-GCN [14] AAAI24 J+B+JM+BM 93.1 97.5 89.2 90.3 -
+BlockGCN [20] CVPR24 J+B+JM+BM 93.1 97.0 90.3 91.5 96.9
+HGCN-based Hyper-GNN [32] TIP21 J+B+JM+BM 89.5 95.7 - - -
+Selective-HCN [33] ICMR22 J+B+JM+BM 90.8 96.6 - - -
+DST-HCN [56] ICME23 J+B+JM+BM 92.3 96.8 88.8 90.7 96.6
+Hyper-GCN (w/o ensemble) - J 91.4 95.5 87.0 88.7 95.5
+Hyper-GCN (4-ensemble) - J+B+JM+BM 93.4 97.1 90.5 91.7 97.1
+TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets.
+For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods.
+J, B, JB, and JM represent the joint, bone, joint motion, and bone motion.
+In addition, for HD-GCN [19], we use the results from 4-streams according to the original papers for a fair comparison.
+The bold font and the underline font represents the best result and 2-nd best result.
 III-E Virtual Connections
@@
 Therefore, we introduce the hyper-joints which are to participate in the hyper-graph convolution as shown in Figure 2 (a).
-The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C^{{}^{\prime}}\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C^{{}^{\prime}}\times T\times V_{h}}$.
+The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C\times T\times V_{h}}$.
 Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data.
@@
 We adopt a cosine matrix $C\in\mathbb{R}^{V_{h}\times V_{h}}$ to measure the differences between hyper-joints, which can be formulated as follows:
-$C=\frac{\text{ReLU}(\mathbf{F_{h}F^{T}_{h}})}{||\mathbf{F_{h}}||^{2}}-\mathbf{I},$ (12)
-where $\mathbf{I}$ represents the identity matrix which is used to emphasise the irrelevance between each learned hyper-joint pair.
-Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer. The calculation can be formulated as:
-$\mathcal{L}_{h}(C)$ $=\frac{\sum^{V_{h}}_{i=1}\sum^{V_{h}}_{j=1}c_{i,j}}{(V_{h}-1)^{2}},\quad c_{i,j}\in C,$ (13)
+$C=\frac{\mathbf{F_{h}F^{T}_{h}}}{||\mathbf{F_{h}}||^{2}},$ (12)
+Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer.
+Since the correlation of hyper-joints themselves cannot be optimised, we manually eliminate this part by minus the $V_{h}$.
+The calculation can be formulated as:
+$\mathcal{L}_{h}(C)$ $=\frac{\sum^{V_{h}}_{i=1}\sum^{V_{h}}_{j=1}ReLU(c_{i,j})-V_{h}}{V_{h}(V_{h}-1)},\quad c_{i,j}\in C,$ (13)
 $\mathcal{L}$ $=\mathcal{L}_{CE}+\frac{1}{L}\sum^{L}_{l=1}L_{h}(C_{l}),$ (14)
-where $V_{h}$ represents the number of hyper-joints we introduced.
-$C_{l}$ represents the cosine matrix of the $l$-th layer.
-Additionally, to preserve the original physical topology, we use the AHC-Module solely to learn the topology between feature $\mathbf{F_{h}}$ and feature $\mathbf{F_{p}}$.
+where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer.
+Additionally, we manually connect the hyper joints to all the physical joints, which optimises the topology from the datasets.
 III-F Entire Architecture
-Hyper-GCN consists of an embedding layer, 9 spatial-temporal convolution layers, a global average pooling, and an FC layer, as shown in Figure 2 (c).
-The embedding layer is utilised to map individual joints into a high-dimensional feature space for subsequent adaptive hyper-graph convolution.
-To preserve the position awareness of the embedded features, we introduce the position embedding in the embedding layer inspired by InfoGCN [12].
-The 9 layers of Hyper-GCN are categorised into 3 stages.
-Each stage consists of 3 layers.
-Each layer consists of Multi-Scale Hyper-graph convolution (MS-HGC) and Multi-Scale Temporal convolution (MS-TC) [13].
-In stage 1, the number of channels is set as 128.
-Then it increases to 256 for stage 2 and stage 3.
-Dense connections are introduced to integrate deep and shallow features in each stage, which can smooth the obtained distribution against the hyper-vertices interaction at each layer.
-Methods Publication Modalities GFLOPs Params (M) NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
-X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
-ST-GCN [18] AAAI 2018 J+B - - 81.5 88.3 70.7 73.2 -
-2s-AGCN [20] CVPR 2019 J+B - - 88.5 95.1 82.5 84.2 -
-DGNN [32] CVPR 2019 J+B+JM+BM - - 89.9 96.1 - - -
-SGN [30] CVPR 2020 J+B+JM+BM - - 89.0 94.5 79.2 81.5 92.5
-DC-GCN+ADG [33] ECCV 2020 J+B+JM+BM 1.83 4.9 90.8 96.6 86.5 88.1 95.3
-DDGCN [31] ECCV 2020 J+B+JM+BM - - 91.1 97.1 - - -
-MS-G3D [42] CVPR 2020 J+B+JM+BM 5.22 2.8 91.5 96.2 86.9 88.4 -
-MST-GCN [34] AAAI 2021 J+B+JM+BM - 12.0 91.5 96.6 87.5 88.8 -
-CTR-GCN [13] ICCV 2021 J+B+JM+BM 1.97 1.5 92.4 96.4 88.9 90.6 96.5
-EfficientGCN-B4 [43] TPAMI 2022 J+B+JM+BM 15.2 2.0 91.7 95.7 88.3 89.1 -
-InfoGCN* [12] CVPR 2022 J+B+J’+B’ 1.84 1.6 92.7 96.9 89.4 90.7 96.6
-FR Head [48] CVPR 2023 J+B+JM+BM - 2.0 92.8 96.8 89.5 90.9 96.8
-HD-GCN* [19] ICCV 2023 J+B+J’+B’ 1.77 1.7 93.0 97.0 89.8 91.2 96.9
-DS-GCN [14] AAAI 2024 J+B+JM+BM - - 93.1 97.5 89.2 90.3 -
-Hyper-GCN (w/o ensemble) - J 1.62 1.0 91.1 95.3 87.0 88.4 94.8
-Hyper-GCN (2-ensemble) - J+B 1.62 1.0 92.9 96.3 89.8 90.9 96.2
-Hyper-GCN (4-ensemble) - J+B+JM+BM 1.62 1.0 93.1 96.7 90.2 91.4 96.6
-TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets.
-For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods.
-J, B, JB, and JM represent the joint, bone, joint motion, and bone motion.
-In addition, for InfoGCN, HD-GCN [12, 19], we use the results from 4-streams according to the original papers for a fair comparison.
-GFLOPs and Params (M) all represent the single-stream.
+Hyper-GCN consists of an embedding layer [12], 9 spatial-temporal convolution layers as shown in Figure 2 (c).
+Each layer consists of the proposed M-HGC and Multi-Scale Temporal convolution (MS-TC) [13].
+9 Layers are categorised into 3 stages with Dense connections.
+The channels in each stage are set to 128, 256 and 512.
 IV Evaluation
-In this section, we conduct extensive experiments on standard benchmarking datasets to evaluate our Hyper-GCN, in the task of skeleton-based human action recognition.
 IV-A Datasets
-NTU-RGB+D 60 NTU-RGB+D 60 [11] is a large dataset widely used in skeleton-based human action recognition.
-It consists of 56,880 different samples, categorised into 60 classes, obtained from the directed performances of 40 different actors.
-There are two 2 evaluation benchmarks.
-(1) Cross-Subject (X-Sub): For the 40 actors, 20 are used for training and 20 for validation.
-(2) Cross-view (X-View): For 3 views, 2 are used for training, and 1 for validation.
-NTU-RGB+D 120 NTU-RGB+D 120 [25] is am extention of NTU 60, which introduce 57,367 new action samples.
-It consists of 114,480 different samples, categorised into 120 classes instead of 60 classes.
-The number of actors increase to 106.
-It also corresponds yo 2 benchmarks.
-(1) Cross-Subject (X-Sub): For the 106 actors, 53 are used for training and 53 for validation.
-(2) Cross-Setup (X-Set): For the 32 setups, the sample with even setup IDs are used for training, and the odd setup IDs for validation.
-Northwestern-UCLA.
-The Northwestern-UCLA (NW-UCLA) dataset [49] contains 1494 video clips, which is categorised into 10 classes.
-It is obtained through 3 cameras with different camera views and is performed by 10 actors.
-As the proposed benchmark, for 3 camera views, 2 camera views are used for training, and the remaining 1 is used for validation.
+NTU-RGB+D 60 & 120 NTU-RGB+D 60 [11] is a large dataset widely used in skeleton-based human action recognition, which is categorised into 60 classes.
+NTU-RGB+D 120 [34] is an extention to 120 classes of NTU 60.
+4 benchmarks recommended by the official are adopted: (1) NTU60-XSub, (2) NTU60-XView, (3) NTU120-XSub, (4) NTU120-Xset.
+Northwestern-UCLA. The Northwestern-UCLA (NW-UCLA) dataset [35] contains 1494 video clips, which is categorised into 10 classes. It contains 3 different camera views and is performed by 10 actors.
 IV-B Implementation Details
 Our training and evaluation stages are on a single RTX 3090 GPU.
-In terms of the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004.
+In the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004.
 Our implementation uses label smooth cross-entropy loss with the Divergence Loss we proposed.
@@
 The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120.
-The training batch size is set to 64 for NTU RGB+D and NTU RGB+D 120, and 16 for Northwestern-UCLA.
 IV-C Comparison with the State-of-the-Art
-Multi-stream ensembles proposed in  [50] have been proven effective by most of the existing state-of-the-art methods. We also use a 4-streams ensemble to evaluate our model.
+Multi-stream ensembles proposed in  [57] have been proven effective by most of the existing state-of-the-art methods.
+We also use a 4-streams ensemble to evaluate our model.
 The detailed results are reported in Table I.
 To the best of our knowledge, a comparison with SOTA shows that our proposed Hyper-GCN achieves the SOTA level on each dataset.
-Especially, we outperform the involved competitors on NTU-RGB+D 120.
 This demonstrates the effectiveness and merits of Hyper-GCN.
 IV-D Ablation Study
-We conduct ablation experiments and report visualisations to analyse the effectiveness of each component of our design.
-All the ablation study is evaluated on the benchmark of NTU-RGB+D 120, using the joint modality.
-The Baseline for comparison is set by replacing the MS-HGC with vanilla GC, and removing the hyper-joints and dense connections in the backbone.
-Figure 4: Top-1 accuracy in selected classes. The $K$-scale represents the $K$-scale hyper-graph in MS-HGC. The data is displayed starting from 75%.
-$K$-scale
-NTU RGB+D 120
-X-Sub (%)
-Baseline - 85.1
-Single scale 3 86.4 ($\uparrow$ 1.3)
-5 86.5 ($\uparrow$ 1.4)
-7 86.5 ($\uparrow$ 1.4)
-9 86.3 ($\uparrow$ 1.2)
-Multi scale [0,2,3,4,5,6,7,8] 86.6 ($\uparrow$ 1.5)
-[2,3,4,5,6,7,8,9] 86.7 ($\uparrow$ 1.6)
-[3,4,5,6,7,8,9,10] 86.6 ($\uparrow$ 1.5)
-TABLE II: Comparison of the hyper-parameter $K$ in terms of adaptive hyper-graph construction.
-[$k_{1}$,$k_{2}$,…,$k_{8}$] represents the $K$-scale hyper-graph of 8 branches in MS-HGC.
-The bold font represents the best results.
-Multi-Scale Hyper-graph: As we mentioned, Multi-Scale Hyper-graph Convolution (MS-HGC) consists of an 8-branch parallel architecture.
-In particular, each branch corresponds to a hyper-graph at a different scale.
-The hyper-graphs at different scales represent the maximum number of hyper-edges each joint is contained in.
+We conduct ablation experiments with visualisations to analyse the effectiveness of our design. All the ablation study is evaluated with the joint modality.
+- NTU RGB+D 120 X-Sub (%)
+K Uniform Non-uniform
+Baseline - 84.7 84.7
+w M-HGC 3 86.4 ($\uparrow$ 1.7) 86.0 ($\uparrow$ 1.3)
+5 86.5 ($\uparrow$ 1.9) 86.2 ($\uparrow$ 1.5)
+7 86.3 ($\uparrow$ 1.6) 86.5 ($\uparrow$ 1.8)
+9 86.0 ($\uparrow$ 1.3) 86.7 ($\uparrow$ 2.2)
+11 85.9 ($\uparrow$ 1.2) 86.4 ($\uparrow$ 1.7)
+TABLE II: Ablation of the hyper-parameter $K$ in A-NHG.
+The Uniform represents K-uniform hyper-graph, in which one hyper-edge contains K joints.
+The non-uniform represents non-uniform hyper-graph, in which one joint is contained by $K$ hyper-edge.
+Figure 4: Visualisation of hyper-graph in selected actions.
+The size of a joint means its weight within the hyper-edge.
+For ease of observation, we visualise the hyper-edges with the larger weight constructed in each sample.
+Hyper-graph Construction: The Non-uniform hyper-graphs with hyper-parameter $K$ represent the maximum number of hyper-edges contained in each joint.
 They reflect the varying degrees of aggregation in the constructed topology.
-The different scales of hyper-graphs are determined by the hyper-parameter $K$ that we define.
-Therefore, to explore how to set the hyper-parameter $K$ for the branches in MS-HGC, we design the ablation experiments to investigate the impact of different scales on the model’s recognition ability, as well as to compare the results by defining a single scale.
+In general, it is common to adopt a uniform hyper-graph with hyper-parameter $K$.
+Therefore, to explore how to set the hyper-parameter $K$ for the adaptive hyper-graph in M-HGC, we design the ablation experiments to investigate the impact of hyper-parameter $K$ on the model’s recognition ability, where both uniform and non-uniform hyper-graphs are considered.
 The results are reported in Table II.
-When $K$ equals 1, each hyper-edge contains only one joint, which essentially eliminates any interaction between joints.
-Therefore, we do not set K to 1.
-Observing the design of the single scale hyper-graph, it becomes clear that increasing $K$ does not always yield better performance.
-A very large $K$ means that the maximum number of hyper-edges each joint is contained in increases, which in turn causes the number of joints in each hyper-edge to grow.
+In both uniform and non-uniform hyper-graphs, a very large $K$ means that more joints are contained in each hyper-edge.
 Even though these ”extra” joints may be optimised for minimal weight, they are likely to act as ”noise”, disrupting the information interaction represented by the hyper-edge.
-Furthermore, compared to a single-scale design, it is evident that the multi-scale design of MS-HGC outperforms the single scale design.
-For the complex and diverse characteristics of human action, multi-scale hyper-graphs in the spatial dimension can better capture features at multiple levels.
-Therefore, we visualise the Top-1 accuracy of selected human action categories as shown in Figure 4.
-This is especially advantageous for recognising actions with varying degrees of motion.
-For instance, for actions like ”Brushing teeth” or ”Salutation”, where only local joints are in motion, the smaller-scale hyper-graphs focus more on the local joints.
-In contrast, for actions like ”Stand up” and ”Jump up”, where the entire body is engaged, a larger-scale hyper-graph can capture information between distant joints.
-This also inspired us to adopt the AHC-Module in different branches to complementarily learn hyper-graph topology at different scales.
-NTU RGB+D 120
-X-Sub (%)
-GFLOPs Parmas (M)
-Baseline 85.1 1.70 1.52
-w/ MS-HGC 86.6 ($\uparrow$ 1.5) 1.59 1.03
-w/ 1 86.7 ($\uparrow$ 1.6) 1.61 1.03
-w/ 2 86.9 ($\uparrow$ 1.8) 1.62 1.04
-w/ 3 87.0 ($\uparrow$ 1.9) 1.62 1.05
-w/ 4 86.6 ($\uparrow$ 1.5) 1.64 1.05
-w/ 5 86.4 ($\uparrow$ 1.3) 1.66 1.06
-TABLE III: The comparison of the configurations of hyper joints.
+As the $K$ decreases, it is difficult for the hyper-edge to represent complex joint combination patterns.
+Therefore, setting $K=5$ in uniform and $K=9$ in Non-uniform achieves the best performance.
+The non-uniform hyper-graph of A-NHG outperforms the uniform hyper-graph, as the combination of skeleton joints in different actions is diverse.
+Furthermore, we visualise the hyper-graph inferred by Hyper-GCN on selected actions in Figure 4.
+The hyper-graph construction focuses on the motion joints that are most relevant to the corresponding action category.
+For “Kicking something” as part (d), the right leg joints and the left hand joint are modeled as a hyper-edge.
+Because in this action, the person’s right hand previously extends, and the right leg swings relatively to the right hand.
+Similarly, in the action ”Stand up” as shown in part (b), most joints of the upper body move relatively to right foot joint, which is not moving in the first half of frames.
+In addition, ”Taking a selfie” and ”Drinking water” are also consistent with the above view as Figure 4 (a), (c).
+NTU RGB+D 120 X-Sub(%)
+w/o $\mathcal{L}_{h}$ w/ $\mathcal{L}_{h}$
+Baseline 84.7 84.7
+w/o M-HGC w/ 1 84.9 ($\uparrow 0.2$) 84.9 ($\uparrow 0.2$)
+w/ 3 84.9 ($\uparrow 0.2$) 85.2 ($\uparrow 0.5$)
+w/ 5 84.7 85.0 ($\uparrow 0.3$)
+w M-HGC w/ 1 86.7 ($\uparrow 2.0$) 86.7 ($\uparrow 2.0$)
+w/ 3 86.6 ($\uparrow 1.8$) 86.9 ($\uparrow 2.2$)
+w/ 5 86.6 ($\uparrow 1.8$) 84.8 ($\uparrow 2.1$)
+TABLE III: The ablations on hyper joints.
 w/ $N$ represents the number of hyper joints in each layer.
-These experiments are conducted in the setting w/ MS-HGC.
-The bold font represents the best results.
-Figure 5: The comparison of the cosine matrices of hyper-joints.
-The matrix elements range from $[-1,1]$.
-This experiment is set w/ 3 hyper-joints.
-Part (a) represents w/o Divergence Loss.
-Part (b) represents w/ Divergence Loss.
-The cosine matrix is calculated by Eqn.
-12.
-Virtual Connections: We conduct the ablation study to analyse the hyper joints with the Divergence Loss.
+w/ $\mathcal{L}_{h}$ represents the training with Divergence Loss.
+The hyper-parameter $K$ of M-HGC is set to $9$.
+Figure 5: Comparison of the cosine matrices of hyper-joints. Part (a) and (b) represents w/o Divergence Loss and w/ Divergence Loss with 3 hyper-joints. It is calculated by Eqn. (12).
+Virtual Connections: We conduct the ablation study to analyse the hyper joints and the Divergence Loss.
 The results are listed in Table III.
@@
 The case of introducing only $3$ hyper joints achieves the best performance.
-As our hyper joints are learned from a large amount of data, too many hyper joints can introduce redundant and ambiguous clues, weakening the model capacity.
-Besides, the involvement of hyper joints delivers consistent improvement, while introducing only a marginal increase in terms of model size and complexity.
-In addition, we visualised the cosine matrices in certain layers of Hyper-GCN, as shown in Figure 5.
-Clearly, in the absence of the Divergence Loss, the homogenisation of the learned hyper-joints in several layers is severe.
-This can limit the ability of hyper-joints to represent the generalised features of human actions.
+As our hyper joints are learned from a large amount of data, involving a large number of hyper joints can introduce redundant and ambiguous clues, degrading the performance.
+In addition, we visualise the cosine matrices of Hyper-GCN, as shown in Figure 5.
+Clearly, in the absence of the Divergence Loss, the homogenisation of the learned hyper-joints is severe.
+This impedes the ability of hyper-joints to represent the generalised features of human actions.
 The observation further validates the effectiveness of the Divergence Loss for hyper-joint optimisation, supporting our viewpoint that generalised features with a certain degree of differentiation are needed to participate in hyper-graph convolution.
-Method S-1 S-2 S-3 MS-HGC H NTU-RGB+D 120
-X-Sub
-Baseline 84.8
-$\checkmark$ 84.8
-$\checkmark$ 84.8
-$\checkmark$ 84.9 ($\uparrow$ 0.1)
-Hyper-GCN $\checkmark$ $\checkmark$ 84.9 ($\uparrow$ 0.1)
-$\checkmark$ $\checkmark$ 85.0 ($\uparrow$ 0.2)
-$\checkmark$ $\checkmark$ $\checkmark$ 85.1 ($\uparrow$ 0.3)
-$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 86.6 ($\uparrow$ 2.0)
-$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 87.0 ($\uparrow$ 2.4)
-TABLE IV: The ablation study of architecture settings. S-$i$ represents adding dense connection in stage $i$. MS-HGC and H represent w/ MS-HGC and w/ 3 hyper-joints.
-Figure 6: t-SNE visualisation of Baseline and Hyper-GCN. Different colours represent the features of different joints.
-Architecture of Backbone: We validate the effectiveness of performing dense connections in the backbone.
-The experimental results are collected in Table IV.
-The dense connections in the experiments are only added within each stage.
-It is observed that dense connections in shallow layers do not improve the performance as much as in deep layers.
-As the layers go deeper, the number of channels gradually increases and the variance of information extracted from each layer becomes steeper.
-This suggests that the deeper stages require dense connections to modulate the variance between features.
-For example, the performance introducing dense connections at stage 1 and stage 2 is close to adding the dense connection only at stage 3.
-To further validate that Hyper-GCN increases the efficiency of information interaction, we perform t-SNE projections on the output features of each layer, as shown in Figure 6.
-Notice that the semantic information represented by the last layer features of Hyper-GCN is very convergent.
-This represents that the semantic information of each joint is adequately conveyed in Hyper-GC.
-The joints with similar semantic can prevent sacrificing the information after global average pooling.
+Figure 6: t-SNE visualisation of Baseline and Hyper-GCN. Different colours represent different joints. Part (a) and (b) reprsent the output features in Baseline and Hyper-GCN.
+Effectiveness of Hyper-GCN: To further validate that Hyper-GCN increases the efficiency of information interaction, we perform t-SNE projections on the output features in last layer between Baseline and Hyper-GCN, as shown in Figure 6 for comparison.
+Notice that the semantic information represented by the last Hyper-GCN layer is very convergent.
+This represents that the semantic information of each joint is adequately conveyed in M-HGC.
+The joints with similar semantics can prevent sacrificing the information after global average pooling.
+Methods Publications NTU RGB+D 120 GFLOPs Params
+X-Sub(%) X-Set(%)
+CTR-GCN[13] ICCV21 88.9 90.4 1.97 1.46
+InfoGCN[12] CVPR23 89.4 90.4 1.70 1.52
+FR-Head[55] CVPR23 89.5 90.9 2.0 -
+BlockGCN[20] CVPR24 90.3 91.5 1.63 1.31
+Hyper-GCN (-) - 90.4 91.5 1.63 1.06
+DSTA-Net[25] ACCV20 86.6 89.0 16.18 3.45
+IIP-Transformer[26] BigData23 89.9 90.9 7.20 2.90
+SkateFormer[44] ECCV24 89.8 91.4 3.62 2.03
+Hyper-GCN - 90.5 91.7 2.88 2.34
+TABLE IV: The comparison of the performance and complexity. Hyper-GCN (-) represents the lightweight version.
+IV-E Comparison of Complexity
+Increasing the number of channels provides considerable improvement, which also increases the complexity to some extent.
+Given that advanced lightweight GCN-based methods are between 1M and 2M.
+For a fair comparison, we set channels in the last 3 layers to 256 to reduce the parameters to 1.06M, which is lighter weight than all involved counterparts.
+The performance of our lightweight version compared with GCN-based methods is reported in Table IV.
+In fact, on NTU RGB+D 120, which is the most challenge dataset, the performance of Hyper-GCN is degraded, but still achieves the SOTA level with the minimal parameter volume.
+In addition, on a comparable basis with the parameters of Transformer-based methods, our Hyper-GCN also outperform the SOTA.
 V Conclusion
-In this paper, we propose a Hyper-graph Convolutional Network for skeleton-based human action recognition.
-To exploit the implicit topology of multivariate synergy between joints, we introduce Multi-Scale Hyper-graph Convolution and virtual connections.
-We also introduce dense connections to fuse shallow and deep features.
+In this paper, we propose an Adaptive Hyper-graph Convolutional Network for skeleton-based human action recognition.
+To exploit the implicit topology of multivariate synergy between joints, we introduce the Adaptive Non-uniform Hyper-graph, the Multi-head Hyper-graph Convolution and virtual connections.
 We carry out experiments on the dataset NTU-RGB+D 60 & 120, NW-UCLA to validate the effectiveness of Hyper-GCN.
@@
 To the best of our knowledge, Hyper-GCN achieves the SOTA performance on 3 public datasets.
-The involvement of hyper-graph modelling indeed extends existing GCN-based action recognition paradigms.
+The involvement of adaptive non-uniform hyper-graph modeling indeed extends existing GCN-based action recognition paradigms.

### v3

@@
 The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods.
-Specifically, we outperform the existing solutions on NTU-120, achieving 90.5% and 91.7% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
-Index Terms:
-Action recognition, Video Understanding, Computer Vision, Graph Learning.
-I Introduction
-Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR, and AR [1, 2, 3, 4, 5, 6, 7].
-In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms.
+The code is available at https://github.com/6UOOON9/Hyper-GCN.
+1 Introduction
+Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR and AR [28, 29, 16, 30, 49, 49, 42, 21].
+In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms [35, 26, 34, 43].
 Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension, conveying human action with higher efficiency.
@@
 Figure 2: The framework of our Hyper-GCN.
-Part (a) represents the Multi-Scale Hyper-GC (MS-HGC) module, where channels are divided into 8 branches.
+Part (a) represents the Multi-head Hyper-graph Convolution (M-HGC) module.
 Part (b) represents the process of constructing an adaptive hyper-graph.
 Part (c) represents the architecture of Hyper-GCN.
+$\mathbf{F_{p}}$ represents the features of real joints.
 $\mathbf{F}_{h}$ represents the hyper-joints which are learnable parameters.
-$\hat{A}_{p}$ represents the physical topology.
-$HA$ represents The fused topology of the adaptive hyper-graph and the human physical structure.
-II Related Work
-II-A Graph Topology for Action Recognition
+$\hat{A}$ represents the physical topology.
+2 Related Work
+2.1 Graph Topology for Action Recognition
 A graph can represent a human skeleton, preserving the joint relationships via predefined edges.
@@
 Though hyper-graph is considered to construct topology using multivariate joint relationships, current [58, 13] solutions manually set the hyper-graphs, which greatly relies on human experience, sacrificing the adaptability of graph learning.
-II-B Feature Configurations for Action Recognition
+2.2 Feature Configurations for Action Recognition
 It has been observed that the input features play the most essential role in delivering high recognition accuracy [5, 27, 41, 1, 15, 47, 31, 33] in GCNs and Transformers.
@@
 Given a fixed representative capacity (number of channels), we believe it is necessary to involve additional virtual joints to balance the pressure of storing local and global information.
-III Approach
-III-A Preliminaries
+3 Approach
+3.1 Preliminaries
 In general, the input features in graph convolution are represented by $\mathbf{{F}_{in}}\in\mathbb{R}^{C\times T\times V}$, where $C$ represents the number of channels in the feature maps.
@@
 $\sigma$ denotes the non-linear activation function ReLU.
-III-B Hyper-graph
+3.2 Hyper-graph
 Here, we use $\mathcal{G_{H}}=(\mathcal{V_{H}},\mathcal{E_{H}},\mathcal{W_{H}})$ to define the spatial hyper-graph [12] with human skeleton.
@@
 $N$ represents the number of joints and $M$ represents the number of hyper-edges.
-Given $v\in\mathcal{V_{H}}$ and $e\in\mathcal{E_{H}}$, the values of incidence matrix can be determined by:
+Given $v\in\mathcal{V_{H}}$ and $e\in\mathcal{E_{H}}$, the values of incidence matrix be determined by:
 $h(e,v)=\begin{cases}1,&v\in e,\\ 0,&v\notin e\end{cases}.$ (3)
@@
 where we use a diagonal matrix $D_{e}\in\mathbb{R}^{M\times M}$ to represent the degree of hyper-edges. Therefore, the normalisation of a hyper-graph is formulated as follows:
-$\hat{H}=D_{v}^{-\frac{1}{2}}HWD_{e}^{-1}H^{T}D_{v}^{-\frac{1}{2}},$ (6)
+$\hat{H}=D_{v}^{-1}HWD_{e}^{-1}H^{T},$ (6)
 where $\hat{H}\in\mathbb{R}^{N\times N}$ represents the normalised incidence matrix for hyper-graph convolution.
@@
 (d) represents the final state.
-III-C Adaptive Non-uniform Hyper-graph
+3.3 Adaptive Non-uniform Hyper-graph
 Constructing the hyper-graph incidence matrix is critical to understanding human actions. As shown in Figure 2 (b), we design an Adaptive Non-uniform Hyper-graph (A-NHG) Construction.
-Given the general feature $\mathbf{X}\in\mathbb{R}^{N\times C}$ of spatial skeleton joints, $N$ represent the number of joints, $C$ represent the channels.
+Given the general features $\mathbf{X}\in\mathbb{R}^{N\times C}$ of spatial skeleton joints, $N$ represent the number of joints, $C$ represent the channels.
 Our A-NHG construct the hyper-edge with each joint as the center of mass, and a total of $N$ hyper-edges are constructed.
@@
 Given $v_{i},v_{j}\in\mathcal{V_{H}}$, each element $m_{i,j}$ in the distance matrix $M\in\mathbb{R}^{N\times N}$ can be obtained as follows:
-$m_{i,j}=m_{j,i}=||v_{i}-v_{j}||^{2}_{2}.$ (7)
+$m_{i,j}=m_{j,i}=||v_{i}-v_{j}||_{2}.$ (7)
 In principle, we need to transform the distance matrix into a probability incidence matrix, which determines whether the specific joints are contained with the same hyper-edges.
@@
 Based on $set_{i}$, $h_{i,j}$ in the incidence matrix $H\in\mathbb{R}^{N\times N}$, can be calculated as follows:
-$h_{i,j}$ =$=\begin{cases}\frac{\text{exp}(-m_{i,j})}{\sum_{k\in set_{i}}{\text{exp}(-m_{i,k})}},&j\in set_{i},\\ 0,&j\notin set_{i}\end{cases}$ (8)
+$h_{i,j}$ =$=\begin{cases}\frac{\text{exp}(-m_{i,j})}{\sum_{k\in set_{i}}{\text{exp}(-m_{i,k})}},&j\in set_{i},\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0,&j\notin set_{i}.\end{cases}$ (8)
 In contrast to uniform hyper-graph (every hyper-edge contains the same number of joints), our approach limits the maximum number of every joint that can be contained by hyper-edges.
@@
 The entire process for example with $K$ = 3 is illustrated in the Figure 3.
-III-D Multi-head Hyper-graph Convolution
+3.4 Multi-head Hyper-graph Convolution
 To accommodate the semantic information reflected by different group channels, we further propose the Multi-head Hyper-graph Convolution (M-HGC), as shown in Figure 2 (a).
 We use separate branches to independently perform hyper-graph convolution on the topologies represented by the multi-head hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN.
-For instance, the feature $\mathbf{F_{in}\in\mathbb{R}^{C_{in}\times T\times V}}$ is divided into $8$ separate branches along the channel dimension, which are processed by $8$ M-HGC parallelly, delivering $8$ separate hyper-graphs.
+For instance, the features $\mathbf{F_{in}}\in\mathbb{R}^{C_{\text{in}}\times T\times V}$ are divided into 888 separate branches along the channel dimension, which are processed by 888 M-HGC parallelly, delivering 888 separate hyper-graphs.
 Next, we perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction.
 This can efficiently decouple the temporal and spatial clues.
-After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}_{in}}\in\mathbb{R}^{C_{in}\times V}$.
+After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}_{in}}\in\mathbb{R}^{C_{\text{in}}\times V}$.
 The hyper-graph is depicted by the incidence matrix and weight matrix.
@@
 After that, we introduce the A-NHG to obtain the incidence matrix.
-For the weight matrix, we adopt the Channel Attention [54].
+For the weight matrix, we adopt MLP to measure each hyper-edge.
 The hyper-graph obtained operation can be formulated as follows:
-$H$ $=\uplus^{8}_{i=1}\text{A-NHG}(\mathbf{\bar{F}^{k}_{in}}),$ (9)
-$W$ $=\sigma^{2}(\Psi^{2}(\uplus^{8}_{i=1}\sigma^{1}(\Psi^{1}_{i}(\mathbf{\bar{F}^{k}_{in}})))),$ (10)
+$H$ $=\uplus^{8}_{k=1}\text{A-NHG}(\mathbf{\bar{F}^{\mathnormal{k}}_{in}}),$ (9)
+$W$ $=\sigma_{2}(\Psi_{2}(\uplus^{8}_{k=1}\sigma_{1}(\Psi^{k}_{1}(\mathbf{\bar{F}^{\mathnormal{k}}_{in}})))),$ (10)
 […]
@@
 The operation of M-HGC can be formulated as follows:
-$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha\cdot\hat{H_{i}})\mathbf{F^{k}_{in}}P_{i}},$ (11)
-$\hat{A}$ and $\hat{H}$ denote the normalized physical adjacency matrix and normalized incidence matrix.
+$\mathbf{F_{out}}$ $=\uplus^{8}_{k=1}{(\hat{A_{k}}+\alpha\cdot\hat{H_{k}})\mathbf{F^{\mathnormal{k}}_{in}}P_{k}},$ (11)
+where $\hat{A}$ and $\hat{H}$ denote the normalized physical adjacency matrix and normalized incidence matrix.
 $P$ is the learnable weight parameter for feature transform.
 $\alpha$ is a learnable parameter for the topology fusion of each head.
-Category Methods Publication Modalities NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
-X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
-GCN-based MST-GCN [40] AAAI21 J+B+JM+BM 91.5 96.6 87.5 88.8 -
-CTR-GCN [13] ICCV21 J+B+JM+BM 92.4 96.4 88.9 90.6 96.5
-EfficientGCN-B4 [49] TPAMI22 J+B+JM+BM 91.7 95.7 88.3 89.1 -
-InfoGCN [12] CVPR22 J+B+JM+BM 92.7 96.9 89.4 90.7 96.6
-FR Head [55] CVPR23 J+B+JM+BM 92.8 96.8 89.5 90.9 96.8
-HD-GCN* [19] ICCV23 J+B+J’+B’ 93.0 97.0 89.8 91.2 96.9
-DS-GCN [14] AAAI24 J+B+JM+BM 93.1 97.5 89.2 90.3 -
-BlockGCN [20] CVPR24 J+B+JM+BM 93.1 97.0 90.3 91.5 96.9
-HGCN-based Hyper-GNN [32] TIP21 J+B+JM+BM 89.5 95.7 - - -
-Selective-HCN [33] ICMR22 J+B+JM+BM 90.8 96.6 - - -
-DST-HCN [56] ICME23 J+B+JM+BM 92.3 96.8 88.8 90.7 96.6
-Hyper-GCN (w/o ensemble) - J 91.4 95.5 87.0 88.7 95.5
-Hyper-GCN (4-ensemble) - J+B+JM+BM 93.4 97.1 90.5 91.7 97.1
-TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets.
-For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods.
-J, B, JB, and JM represent the joint, bone, joint motion, and bone motion.
-In addition, for HD-GCN [19], we use the results from 4-streams according to the original papers for a fair comparison.
-The bold font and the underline font represents the best result and 2-nd best result.
-III-E Virtual Connections
+3.5 Virtual Connections
 It is worth emphasising that incorporating learnable joints among different samples is essential for enhancing the model capacity, as these learnable joints capture generalised features of human actions.
@@
 Therefore, we introduce the hyper-joints which are to participate in the hyper-graph convolution as shown in Figure 2 (a).
-The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C\times T\times V_{h}}$.
+The shape of hyper-joints is consistent with the features of physical joints $\mathbf{F_{p}}\in\mathbb{R}^{C\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C\times T\times V_{h}}$ ($V_{h}$ represents the number of hyper-joints).
+To ensure alignment in the temporal dimension, the hyper-joints are shared across each frame.
 Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data.
@@
 We adopt a cosine matrix $C\in\mathbb{R}^{V_{h}\times V_{h}}$ to measure the differences between hyper-joints, which can be formulated as follows:
-$C=\frac{\mathbf{F_{h}F^{T}_{h}}}{||\mathbf{F_{h}}||^{2}},$ (12)
+$C=\frac{\mathbf{F_{h}F^{\mathnormal{T}}_{h}}}{||\mathbf{F_{h}}||^{2}},$ (12)
 Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer.
 Since the correlation of hyper-joints themselves cannot be optimised, we manually eliminate this part by minus the $V_{h}$.
-The calculation can be formulated as:
+The loss calculation can be formulated as:
 $\mathcal{L}_{h}(C)$ $=\frac{\sum^{V_{h}}_{i=1}\sum^{V_{h}}_{j=1}ReLU(c_{i,j})-V_{h}}{V_{h}(V_{h}-1)},\quad c_{i,j}\in C,$ (13)
 $\mathcal{L}$ $=\mathcal{L}_{CE}+\frac{1}{L}\sum^{L}_{l=1}L_{h}(C_{l}),$ (14)
-where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer.
+where $V_{h}$ represents the number of hyper-joints we introduced.
+$C_{l}$ represents the cosine matrix of the $l$-th layer.
+$\mathcal{L}_{\text{CE}}$ represents the cross-entropy loss.
 Additionally, we manually connect the hyper joints to all the physical joints, which optimises the topology from the datasets.
-III-F Entire Architecture
+3.6 Entire Architecture
 Hyper-GCN consists of an embedding layer [6], 9 spatial-temporal convolution layers as shown in Figure 2 (c).
-Each layer consists of the proposed M-HGC and Multi-Scale Temporal convolution (MS-TC) [13].
-9 Layers are categorised into 3 stages with Dense connections.
-The channels in each stage are set to 128, 256 and 512.
-IV Evaluation
-IV-A Datasets
+Each layer consists of the proposed M-HGC and Multi Scale Temporal convolution (MS-TC) [3].
+9 Layers are categorised into 3 stages.
+In each stage, we introduce the dense connections to integrate deep and shallow features in each stage.
+Additionally, to further validate the potential of Hyper-GCN, we propose a base version and a large version.
+The channels in each stage are set to 128, 256, 256 for the base version and 128, 256, 512 for the large version.
+Category Methods Modalities Params (M) GFLOPs NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
+X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
+GCN ST-GCN [52] J+B - - 81.5 88.3 70.7 73.2 -
+2s-AGCN [36] J+B - - 88.5 95.1 82.5 84.2 -
+DC-GCN+ADG [4] J+B+JM+BM 4.9 1.83 90.8 96.6 86.5 88.1 95.3
+MS-G3D [27] J+B+JM+BM 2.8 5.22 91.5 96.2 86.9 88.4 -
+MST-GCN [11] J+B+JM+BM 12.0 - 91.5 96.6 87.5 88.8 -
+CTR-GCN [3] J+B+JM+BM 1.5 1.97 92.4 96.4 88.9 90.6 96.5
+EfficientGCN-B4 [41] J+B+JM+BM 2.0 15.20 91.7 95.7 88.3 89.1 -
+InfoGCN [6] J+B+JM+BM 1.6 1.84 92.7 96.9 89.4 90.7 96.6
+FR Head [56] J+B+JM+BM 2.0 - 92.8 96.8 89.5 90.9 96.8
+HD-GCN* [23] J+B+J’+B’ 1.7 1.77 93.0 97.0 89.8 91.2 96.9
+DS-GCN [51] J+B+JM+BM - - 93.1 97.5 89.2 90.3 -
+BlockGCN [57] J+B+JM+BM 1.3 1.63 93.1 97.0 90.3 91.5 96.9
+Transformer DSTA-Net [38] J+B+JM+BM 3.5 16.18 89.5 95.7 86.6 89.0 -
+IIP-Transformer [47] J+B+JM+BM 2.9 7.20 89.5 95.7 89.9 90.9 -
+SkateFormer [8] J+B+JM+BM 2.0 3.62 93.5 97.8 89.8 91.4 98.3
+HGCN Hyper-GNN [13] J+B+JM+BM - - 89.5 95.7 - - -
+Selective-HCN [58] J+B+JM+BM - - 90.8 96.6 - - -
+DST-HCN [48] J+B+JM+BM 3.5 2.93 92.3 96.8 88.8 90.7 96.6
+Ours (B) J+B+JM+BM 1.1 1.63 93.3 97.4 90.5 91.7 97.2
+Ours (L) J+B+JM+BM 2.3 2.88 93.7 97.8 90.9 92.0 97.6
+Table 1: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets.
+For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods.
+J, B, JB, and JM represent the joint, bone, joint motion, and bone motion.
+In addition, for HD-GCN [23], we use the results from 4-streams according to the original papers for a fair comparison.
+The bold font represents the best result and underline font represents the 2-nd best result.
+4 Evaluation
+4.1 Datasets
 NTU-RGB+D 60 & 120 NTU-RGB+D 60 [35] is a large dataset widely used in skeleton-based human action recognition, which is categorised into 60 classes.
@@
 Northwestern-UCLA. The Northwestern-UCLA (NW-UCLA) dataset [46] contains 1494 video clips, which is categorised into 10 classes. It contains 3 different camera views and is performed by 10 actors.
-IV-B Implementation Details
-Our training and evaluation stages are on a single RTX 3090 GPU.
+4.2 Implementation Details
+Our training and evaluation stages are on a single GPU RTX 3090.
 In the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004.
@@
 The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120.
-IV-C Comparison with the State-of-the-Art
-Multi-stream ensembles proposed in  [57] have been proven effective by most of the existing state-of-the-art methods.
-We also use a 4-streams ensemble to evaluate our model.
-The detailed results are reported in Table I.
-To the best of our knowledge, a comparison with SOTA shows that our proposed Hyper-GCN achieves the SOTA level on each dataset.
+4.3 Comparison with the State-of-the-Art
+Multi-stream ensemble proposed in  [39] have been proven effective by most of the existing state-of-the-art methods.
+We also use a 4-stream ensemble to evaluate the base version and the large version of Hyper-GCN.
+The detailed results are reported in Table 1.
+In summary, the base version of Hyper-GCN comprehensively outperforms all GCN-based and HGCN-based (hyper-graph utilised) SOTA and surpasses the Transformer-based SOTA on the NTU120, with a most lightweight design.
+Moreover, when scaled to match the parameters of the lightest Transformer-based method, the large version of Hyper-GCN achieves 1-st place in 4 benchmarks and 2-nd place in 1 benchmark.
 This demonstrates the effectiveness and merits of Hyper-GCN.
-IV-D Ablation Study
-We conduct ablation experiments with visualisations to analyse the effectiveness of our design. All the ablation study is evaluated with the joint modality.
+4.4 Ablation Study
+We conduct ablation experiments with visualisations to analyse the effectiveness of our design. redAll the ablation study is evaluated with the joint modality of base version.
 NTU RGB+D 120 X-Sub (%)
@@
 11 85.9 ($\uparrow$ 1.2) 86.4 ($\uparrow$ 1.7)
-TABLE II: Ablation of the hyper-parameter $K$ in A-NHG.
+Table 2: Ablation of the hyper-parameter $K$ in A-NHG.
 The Uniform represents K-uniform hyper-graph, in which one hyper-edge contains K joints.
 The non-uniform represents non-uniform hyper-graph, in which one joint is contained by $K$ hyper-edge.
-Figure 4: Visualisation of hyper-graph in selected actions.
-The size of a joint means its weight within the hyper-edge.
-For ease of observation, we visualise the hyper-edges with the larger weight constructed in each sample.
+Figure 4: Visualisation of hyper-graph in selected actions. For ease of observation, the joints contained in the selected hyper-edge are highlighted in red.
 Hyper-graph Construction: The Non-uniform hyper-graphs with hyper-parameter $K$ represent the maximum number of hyper-edges contained in each joint.
@@
 Therefore, to explore how to set the hyper-parameter $K$ for the adaptive hyper-graph in M-HGC, we design the ablation experiments to investigate the impact of hyper-parameter $K$ on the model’s recognition ability, where both uniform and non-uniform hyper-graphs are considered.
-The results are reported in Table II.
+The results are reported in Table 2.
 In both uniform and non-uniform hyper-graphs, a very large $K$ means that more joints are contained in each hyper-edge.
@@
 Furthermore, we visualise the hyper-graph inferred by Hyper-GCN on selected actions in Figure 4.
-The hyper-graph construction focuses on the motion joints that are most relevant to the corresponding action category.
-For “Kicking something” as part (d), the right leg joints and the left hand joint are modeled as a hyper-edge.
-Because in this action, the person’s right hand previously extends, and the right leg swings relatively to the right hand.
-Similarly, in the action ”Stand up” as shown in part (b), most joints of the upper body move relatively to right foot joint, which is not moving in the first half of frames.
-In addition, ”Taking a selfie” and ”Drinking water” are also consistent with the above view as Figure 4 (a), (c).
+The hyper-graph construction focuses on the joints that are most relevant to the corresponding action category.
+For “Kicking something” as part (d), the left leg joints and the right hand joints are modelled as a hyper-edge.
+Because in this action, the person’s right hand is previously extended forward, and the left leg swings relatively to the right hand.
+Similarly, in the action ”Stand up” as shown in part (b), the hand and foot joints serve as fixed joints, while the remaining body joints move relative to them.
+In addition, ”Taking a selfie” and ”Drinking water” both focus on the main joints of movement with the above view as Figure 4 (a), (c).
 NTU RGB+D 120 X-Sub (%)
@@
 Baseline 84.7 84.7
-w/o M-HGC w/ 1 84.9 ($\uparrow 0.2$) 84.9 ($\uparrow 0.2$)
-w/ 3 84.9 ($\uparrow 0.2$) 85.2 ($\uparrow 0.5$)
-w/ 5 84.7 85.0 ($\uparrow 0.3$)
-w M-HGC w/ 1 86.7 ($\uparrow 2.0$) 86.7 ($\uparrow 2.0$)
-w/ 3 86.6 ($\uparrow 1.8$) 86.9 ($\uparrow 2.2$)
-w/ 5 86.6 ($\uparrow 1.8$) 84.8 ($\uparrow 2.1$)
-TABLE III: The ablations on hyper joints.
+w/o M-HGC w/ 1 84.9 (0.2$\uparrow$ 0.2↑ 0.2) 84.9 (0.2$\uparrow$ 0.2↑ 0.2)
+w/ 3 84.9 (0.2$\uparrow$ 0.2↑ 0.2) 85.2 (0.5$\uparrow$ 0.5↑ 0.5)
+w/ 5 84.7 85.0 (0.3$\uparrow$ 0.3↑ 0.3)
+w M-HGC w/ 1 86.7 (2.0$\uparrow$ 2.0↑ 2.0) 86.7 (2.0$\uparrow$ 2.0↑ 2.0)
+w/ 3 86.6 (1.8$\uparrow$ 1.8↑ 1.8) 86.9 (2.2$\uparrow$ 2.2↑ 2.2)
+w/ 5 86.6 (1.8$\uparrow$ 1.8↑ 1.8) 86.8 (2.1$\uparrow$ 2.1↑ 2.1)
+Table 3: Ablations on hyper joints.
 w/ $N$ represents the number of hyper joints in each layer.
 w/ $\mathcal{L}_{h}$ represents the training with Divergence Loss.
-The hyper-parameter $K$ of M-HGC is set to $9$.
+The hyper-parameter $K$ of M-HGC is set to 999.
 Figure 5: Comparison of the cosine matrices of hyper-joints. Part (a) and (b) represents w/o Divergence Loss and w/ Divergence Loss with 3 hyper-joints. It is calculated by Eqn. (12).
 Virtual Connections: We conduct the ablation study to analyse the hyper joints and the Divergence Loss.
-The results are listed in Table III.
+The results are listed in Table 3.
 By comparison, the performance is not monotonically increasing with the number of hyper joints.
-The case of introducing only $3$ hyper joints achieves the best performance.
+The case of introducing only 333 hyper joints achieves the best performance.
 As our hyper joints are learned from a large amount of data, involving a large number of hyper joints can introduce redundant and ambiguous clues, degrading the performance.
@@
 The joints with similar semantics can prevent sacrificing the information after global average pooling.
-Methods Publications NTU RGB+D 120 GFLOPs Params
-X-Sub(%) X-Set(%)
-CTR-GCN[13] ICCV21 88.9 90.4 1.97 1.46
-InfoGCN[12] CVPR23 89.4 90.4 1.70 1.52
-FR-Head[55] CVPR23 89.5 90.9 2.0 -
-BlockGCN[20] CVPR24 90.3 91.5 1.63 1.31
-Hyper-GCN (-) - 90.4 91.5 1.63 1.06
-DSTA-Net[25] ACCV20 86.6 89.0 16.18 3.45
-IIP-Transformer[26] BigData23 89.9 90.9 7.20 2.90
-SkateFormer[44] ECCV24 89.8 91.4 3.62 2.03
-Hyper-GCN - 90.5 91.7 2.88 2.34
-TABLE IV: The comparison of the performance and complexity. Hyper-GCN (-) represents the lightweight version.
-IV-E Comparison of Complexity
-Increasing the number of channels provides considerable improvement, which also increases the complexity to some extent.
-Given that advanced lightweight GCN-based methods are between 1M and 2M.
-For a fair comparison, we set channels in the last 3 layers to 256 to reduce the parameters to 1.06M, which is lighter weight than all involved counterparts.
-The performance of our lightweight version compared with GCN-based methods is reported in Table IV.
-In fact, on NTU RGB+D 120, which is the most challenge dataset, the performance of Hyper-GCN is degraded, but still achieves the SOTA level with the minimal parameter volume.
-In addition, on a comparable basis with the parameters of Transformer-based methods, our Hyper-GCN also outperform the SOTA.
-V Conclusion
+5 Conclusion
 In this paper, we propose an Adaptive Hyper-graph Convolutional Network for skeleton-based human action recognition.
@@
 To the best of our knowledge, Hyper-GCN achieves the SOTA performance on 3 public datasets.
-The involvement of adaptive non-uniform hyper-graph modeling indeed extends existing GCN-based action recognition paradigms.
+The involvement of adaptive non-uniform hyper-graph modelling indeed extends existing GCN-based action recognition paradigms.
+Acknowledgment This work was supported in part by the National Natural Science Foundation of China (62020106012, 62106089, 62332008, 62336004), the Engineering and Physical Sciences Research Council (EPSRC), U.K.
+(Grants EP/V002856/1, and EP/T022205/1), and the Fundamental Research Funds for the Central Universities (JUSRP202504007).